### v3

# Bank–Laine functions with preassigned number of zeros

Abstract.
A Bank–Laine function $E$ is written as $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of the second order differential equation $f^{\prime\prime}+Af=0$, where $A$ is an entire function. In this paper, we first complete the construction of Bank–Laine functions by Bergweiler and Eremenko and then show the existence of entire functions $A$ for which the associated Bank–Laine functions $E=f_{1}f_{2}$ have preassigned number of zeros of three types:
  • (1)
    for every positive integer $n\in\mathbb{N}$ and every two numbers $\lambda_{1},\lambda_{2}\in[0,n]$ such that $\lambda_{1}\leq\lambda_{2}$, there exists an entire function $A$ of order $n$ such that $E=f_{1}f_{2}$ satisfies $\lambda(f_{1})=\lambda_{1}$, $\lambda(f_{2})=\lambda_{2}$ and $\lambda(E)=\lambda_{2}\leq\rho(E)=n$;
  • (2)
    for every number $\rho\in(1/2,\infty)$ and every number $\lambda\in[0,\infty)$, there exists an entire function $A$ of order $\rho$ such that $E=f_{1}f_{2}$ satisfies $\lambda(f_{1})=\lambda$ and $\lambda(f_{2})=\infty$ and $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other $f_{3}$ linearly independent from $f_{1}$;
  • (3)
    for every positive integer $n\in\mathbb{N}$ and every two numbers $\lambda_{1}\in[0,n]$ and $\lambda_{2}\in[0,\infty]$, there exists an entire function $A$ of order $n$ such that $E=f_{1}f_{2}$ satisfies $\lambda(f_{1})=\lambda_{1}$, $\lambda(f_{2})=\lambda_{2}$ and $\lambda(E)=\max\{\lambda_{1},\lambda_{2}\}\leq\rho(E)=\infty$ and $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other $f_{3}$ linearly independent from $f_{1}$.
The constructions for the three types of Bank–Laine functions require new developments of the method of quasiconformal surgery by Bergweiler and Eremenko.
Key words and phrases:
Linear differential equation; Bank–Laine function; Bank–Laine conjecture; Quasiregular function; Quasiconformal surgery
2010 Mathematics Subject Classification:
Primary 34A20; Secondary 30D15
The author is supported by the Fundamental Research Funds for the Central Universities (FRF-TP-19-055A1) and a Project supported by the National Natural Science Foundation of China (12301091)
1. Introduction
In the last 1980s, Bank and Laine [2, 3] initiated the study on the complex oscillation of the second order differential equation
(1.1) $f^{\prime\prime}+Af=0,$
where $A$ is an entire function, under the framework of Nevanlinna theory. See [10] for the standard notation and fundamental results of Nevanlinna theory. Recall that, for a nonconstant entire function $f$, the order $\rho(f)$ of $f$ is defined as
$\rho(f)=\limsup_{r\to\infty}\frac{\log\log|f(z)|}{\log|z|}$
and the exponent of convergence $\lambda(f)$ of zeros of $f$ is defined as
$\lambda(f)=\inf\left\{\lambda:\sum_{z\not=0:f(z)=0}m(z)|z|^{-\lambda}<\infty\right\}=\limsup_{r\to\infty}\frac{\log n(r,0,f)}{\log r},$
where $m(z)$ is the multiplicity of the zero $z$ and $n(r,0,f)=\sum_{|z|\leq r}m(z)$ denotes the number of zeros of $f$ in the disc $\{z:|z|\leq r\}$. It is elementary that all nontrivial solutions of equation (1.1) are entire functions. When $A$ is a transcendental entire function of finite order, an application of the lemma on the logarithmic derivative [10, Section 2.2] easily yields that all nontrivial solutions of (1.1) satisfy $\rho(f)=\infty$. However, there are examples of $A$ (say $A(z)=e^{z}-q^{2}/16$ for an odd integer $q$) such that the product $E=f_{1}f_{2}$ of two linearly independent solutions of (1.1) satisfies $\lambda(E)<\infty$ or even $\lambda(E)=0$ [12, Theorem 5.22]. Generally, we have the following
Theorem 1.1.
Let $E$ be the product of two linearly independent solutions $f_{1}$ and $f_{2}$ of the differential equation (1.1). Then
  • (1)
    if $\rho(A)\not\in\mathbb{N}$, then $\lambda(E)\geq\rho(A)$;
  • (2)
    if $\rho(A)\leq 1/2$, then $\lambda(E)=\infty$.
These results were proved by Bank and Laine in [2, 3], except for the case $\rho(A)=1/2$ in (2) which was prove by Shen [16] and Rossi [15] independently. Shen [16] and Rossi [15] actually showed that if $11/2\leq\rho(A)<$ 1, then the conclusion $\lambda(E)\geq\rho(A)$ can be improved to satisfy the inequality
(1.2) $\frac{1}{\rho(A)}+\frac{1}{\lambda(E)}\leq 2.$
A problem left by their work–known as the Bank–Laine conjecture–is whether $\lambda(E)=\infty$ whenever $\rho(A)\not\in\mathbb{N}$. This conjecture has attracted much interest; see the surveys [8, 13] and references therein. Bergweiler and Eremenko [4, 5] disproved this conjecture by using the method of quasiconformal surgery to construct, for each $\rho\in(1/2,\infty)$, entire function $A$ such that $\rho(A)=\rho$ and the differential equation (1.1) has two linearly independent solutions $f_{1},f_{2}$ such that $\lambda(E)=\rho(E)<\infty$ and one of $f_{1}$ and $f_{2}$ has no zeros. We refer to [1, 14] for the basic results of quasiconformal mapping theory.
An entire function $E$ is called a Bank–Laine function if $E(z)=0$ implies that $E^{\prime}(z)\in\{-1,1\}$. In particular, $E$ is called special if $E(z)=0$ implies that $E^{\prime}(z)=1$. For two linearly independent solutions $f_{1}$ and $f_{2}$ of (1.1), if the Wronskian determinant $W(f_{1},f_{2})=f_{1}f_{2}^{\prime}-f_{1}^{\prime}f_{2}=1$, then we say that $f_{1}$ and $f_{2}$ are normalized. A Bank–Laine function $E$ can be written as $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of the second order differential equation (1.1); see [12, Proposition 6.4]. Most study on the Bank–Laine conjecture has concentrated on constructing Bank–Laine functions with certain properties. All Bank–Laine functions have order at least $1$ by Theorem 1.1 and the inequality (1.2). Bergweiler and Eremenko’s construction in [4] shows for the first time that there exists a special Bank–Laine function $E$ such that $\rho(E)$ is not an integer order or half-integer order.
For two linearly independent solutions $f_{1}$ and $f_{2}$ of equation (1.1), it is well known that the ratio $F=f_{2}/f_{1}$ satisfies the Schwarzian differential equation
(1.3) $\begin{split}S(F):=\left(\frac{F^{\prime\prime}}{F^{\prime}}\right)^{\prime}-\frac{1}{2}\left(\frac{F^{\prime\prime}}{F^{\prime}}\right)^{2}=\frac{F^{\prime\prime\prime}}{F^{\prime}}-\frac{3}{2}\left(\frac{F^{\prime\prime}}{F^{\prime}}\right)^{2}=2A,\end{split}$
where the expression $S(F)$ is called the Schwarzian derivative of $F$. Then $F$ is a locally univalent meromorphic function and the Schwarzian derivative is invariant under the linear fractional transformation of $F$. This implies a one-to-one correspondence between the equivalence class of locally univalent meromorphic functions and entire function. On the other hand, a normalized pair $(f_{1},f_{2})$ of solutions of (1.1) can be recovered from $F$ by the formulas $f_{1}^{2}=1/F^{\prime}$ and $f_{2}^{2}=F^{2}/F^{\prime}$. The quotient $F/F^{\prime}=f_{1}f_{2}=E$ is then a Bank–Laine function and all Bank–Laine functions arise in this way. Zeros of a Bank–Laine function $E$ are zeros and poles of $F$ respectively, and $E$ is special if and only if $F$ is entire. By gluing different locally univalent meromorphic functions together using quasiconformal surgery, Bergweiler and Eremenko then recover the desired $F$ and also the corresponding coefficient $A$ from equation (1.3). In [19], the present author completed Bergweiler and Eremenko’s construction in [5] and, in particular, prove the following
Theorem 1.2 ([19]).
Let $\lambda\in(1,\infty)$ and $\delta\in[0,1]$. Then there exists a Bank–Laine function $E$ such that $\lambda(E)=\rho(E)=\lambda$ and $E=f_{1}f_{2}$ for two entire functions $f_{1}$ and $f_{2}$ such that $\lambda(f_{1})=\delta\lambda$ and $\lambda(f_{2})=\lambda$.
The construction in [19] have started from a family of locally univalent meromorphic functions $F=f_{2}/f_{1}$ with $f_{1}$ and $f_{2}$ being two normalized solutions such that $\lambda(f_{i})<\infty$, $i=1,2$, of the second order differential equation
$f^{\prime\prime}-\left[\frac{1}{4}e^{2z}+\left(\frac{k_{2}-k_{1}}{2}\right)e^{z}+\frac{(k_{1}+k_{2}+1)^{2}}{4}\right]f=0,$
where $k_{1},k_{2}$ are two nonnegative integers (see also [6, 21, 20, 18]). This equation admits two normalized solutions $f_{1}$ and $f_{2}$ in the form
$\begin{split}f_{1}=\kappa_{1}e^{-\frac{k_{1}+k_{2}+1}{2}z}e^{-\frac{1}{2}e^{z}}=\left(\sum_{i=0}^{k_{1}}a_{i}e^{iz}\right)e^{-\frac{k_{1}+k_{2}+1}{2}z}e^{-\frac{1}{2}e^{z}}\end{split}$
and
$\begin{split}f_{2}=\kappa_{2}e^{-\frac{k_{1}+k_{2}+1}{2}z}e^{\frac{1}{2}e^{z}}=\left(\sum_{j=0}^{k_{2}}b_{j}e^{jz}\right)e^{-\frac{k_{1}+k_{2}+1}{2}z}e^{\frac{1}{2}e^{z}}\end{split}$
with
$\left\{\begin{array}[]{ll}a_{i}=\frac{k_{2}!}{(i+k_{2})!}a_{0},&\quad i=1,\cdots,k_{1},\\ b_{j}=\frac{(-1)^{j}k_{1}!}{(j+k_{1})!}b_{0},&\quad j=1,\cdots,k_{2}\end{array}\right.$
and $a_{0}$ and $b_{0}$ are two constants such that $a_{k_{1}}b_{k_{2}}=\frac{k_{1}!k_{2}!}{[(k_{1}+k_{2})!]^{2}}a_{0}b_{0}=1$. Let
$\begin{split}F=\frac{f_{2}}{f_{1}}=\left(\frac{\sum_{j=0}^{k_{2}}b_{j}e^{jz}}{\sum_{i=0}^{k_{1}}a_{i}e^{iz}}\right)e^{e^{z}}.\end{split}$
We choose $a_{0}=b_{0}$. Let $A_{0}=B_{0}=1$ and $A_{i}=a_{i}/a_{0}$ and $B_{j}=b_{j}/b_{0}$. For every pair $(m,2n)$ of two nonnegative integers $m$ and $n$, we consider the rational function
(1.4) $\begin{split}R_{m,n}(z)=\frac{P_{n}(z)}{Q_{m}(z)}=\frac{\sum_{j=0}^{2n}B_{j}z^{j}}{\sum_{i=0}^{m}A_{i}z^{i}}.\end{split}$
Thus we may write
$\begin{split}h_{m,n}(z)=R_{m,n}(z)e^{z}=\left(\frac{\sum_{j=0}^{2n}B_{j}z^{j}}{\sum_{i=0}^{m}A_{i}z^{i}}\right)e^{z}.\end{split}$
It follows that $A_{m}B_{2n}=\frac{m!(2n)!}{[(m+2n)!]^{2}}=\frac{1}{\binom{m+2n}{m}(m+2n)!}$. Then the meromorphic (entire when $m=0$) function
(1.5) $\begin{split}g_{m,n}(z)=h_{m,n}(e^{z})=\left(\frac{\sum_{j=0}^{2n}B_{j}e^{jz}}{\sum_{i=0}^{m}A_{i}e^{iz}}\right)e^{e^{z}}\end{split}$
satisfies
(1.6) $\begin{split}g^{\prime}_{m,n}(z)=\frac{1}{\binom{m+2n}{m}(m+2n)!}\frac{e^{e^{z}+(m+2n+1)z}}{\left(\sum_{i=0}^{m}A_{i}e^{iz}\right)^{2}}.\end{split}$
Since $A_{j}$ are all positive numbers, $g^{\prime}_{m,n}(z)\not=0$ for all $z\in\mathbb{C}$ and $g_{m,n}$ is increasing on $\mathbb{R}$, and satisfies $g_{m,n}(x)\to 1$ as $x\to-\infty$ as well as $g_{m,n}(x)\to\infty$ as $x\to\infty$. By the recursive formulas for $A_{i}$, we see that if $n\geq 1$, then $A_{1}+A_{2}+\cdots+A_{m}<1$ for all $m\geq 1$, which implies that $g_{m,n}$ has no poles on the imaginary axis. If $n=0$ and $m\geq 1$, it is possible that $g_{m,n}$ has some poles on the imaginary axis. Bergweiler and Eremenko’s glued the functions $g_{m,n}$ in (1.5) with $m=0$ and two or infinitely many different even numbers $2n$. Our construction in [19] considered the generic choice of $m$ and $n$. For the underlying ideas behind the gluing, see [5, Section 2].
To prove Theorem 1.2, for each number $\rho\in(1/2,1)$, we let $\gamma=1/(2\rho-1)$. In [19], the author actually proved that, for every $\delta\in[0,1]$, there exists an entire function $A$ of order $\rho(A)=\rho$ such that the differential equation (1.1) has two linearly independent solutions $f_{1}$ and $f_{2}$ such that $E=f_{1}f_{2}$ satisfies $\lambda(E)=\rho\gamma$ and $f_{1}$ and $f_{2}$ satisfy $\lambda(f_{1})=\delta\rho\gamma$ and $\lambda(f_{2})=\rho\gamma$, respectively. In particular, when $\delta=0$ the function $f_{1}$ can be chosen to have no zeros so that the associated Bank–Laine function $E$ is special. Theorem 1.2 completes the construction in [5, Corollary 1.1]. As pointed out in [19], we can extend the results to the case that $\rho\in(n/2,n)$ for any positive integer $n$ by following exactly the same process as in [5, Section 5], thus completing the construction in [5, Corollary 1.2]; see also the comments in Subsection 3.5 and Subsection 4.3. Unfortunately, the proof of Theorem 1.2 in [19] is not complete because the proof for the two lemmas in [19, Lemma 2.1 and Lemma 2.2] contains gaps in both of the two cases $\gamma=1$ and $\gamma>1$. One of the purposes of this paper is to fill these gaps. The case $\gamma=1$ will be included in Theorem 1.4 below and the corrected versions of the two lemmas will be provided in Section 4.
We note that, when $\delta=1$ and $\gamma>1$, the zeros of the function $f_{1}$ constructed in Theorem 1.2 actually satisfies
(1.7) $c_{2}(\log r)^{-2}r^{\lambda}\leq n(r,0,f_{1})\leq c_{1}(\log r)^{-2}r^{\lambda}$
for two positive constants $c_{1}$ and $c_{2}$ and all large $r$. Thus $f_{1}$ has less number of zeros than $f_{2}$. It seems that the method in [19] does not allow to remove the term $(\log r)^{-2}$ in (1.7). For two positive quantities $S_{1}(r)$ and $S_{2}(r)$, we say that $S_{1}(r)\asymp S_{2}(r)$, provided that there are two positive constants $c_{1}$ and $c_{2}$ such that $S_{1}(r)$ and $S_{2}(r)$ satisfy $c_{2}S_{2}(r)\leq S_{1}(r)\leq c_{1}S_{2}(r)$ for all large $r$. As a complement to Theorem 1.2, we shall first consider gluing two locally univalent meromorphic functions $g_{m,n}$ in (1.5) along the real axis as in [4] and prove the following
Theorem 1.3.
Let $M$ and $N$ be two positive integers such that $M\not=N$. Then there exists an entire function $A$ of order
$\rho=1+\frac{\log^{2}(M/N)}{4\pi^{2}}$
such that the associated Bank–Laine function $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(E)=\rho(E)=\rho$ and the number of zeros of $f_{1}$ and $f_{2}$ satisfy $n(r,0,f_{1})\asymp n(r,0,f_{2})$.
By the proof of Theorem 1.3, we know that the zeros of $f_{1}$ and $f_{2}$ actually satisfy $n(r,0,f_{i})\asymp r^{1/\text{Re}(\mu)}(\log r)$, $i=1,2$. Moreover, if we assume that $M=N$, then the term $\log r$ here can be removed. We remark that we can choose $M$ and $N$ suitably so that $f_{1}$ has no zeros and thus the corresponding Bank–Laine function $E$ is special. We may even choose $M$ and $N$ suitably so that $f_{2}$ has no zeros.
By looking at the results in Theorem 1.1, Theorem 1.2 and Theorem 1.3, the existence of Bank–Laine functions in the case $\rho(A)$ is an integer is still not clear. Langley [11] pointed out, using Lagrange interpolation, we can make Bank–Laine functions $E=Pe^{Q}$ with $P,Q$ being polynomials. However, it is not known if there exist Bank–Laine functions $E$ such that $\lambda(E)=\lambda<\rho(E)=n$ for any preassigned number $\lambda$ and positive integer $n$. Moreover, all the constructions now yield Bank–Laine functions such that $\rho(E)=\rho(A)$, but it is not known if there exist Bank–Laine functions such that $\lambda(E)>\rho(A)$. It seems that the method of quasiconformal surgery in [5, Subsection 5.3] is invalid when dealing with these two cases, as will be explained in Section 3. In this paper we shall describe a new method of quasiconformal surgery for constructions of such kind of Bank–Laine functions. Concerning the first question, we shall prove the following
Theorem 1.4.
Let $n\in\mathbb{N}$ be a positive integer and $\lambda_{1},\lambda_{2}\in[0,n]$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$. Then there exists an entire function $A$ of order $n$ such that the associated Bank–Laine function $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})=\lambda_{1}$, $\lambda(f_{2})=\lambda_{2}$ and $\lambda(E)=\lambda_{2}\leq\rho(E)=n$.
In Theorem 1.4, the number of zeros of $f_{1}$ actually satisfies $n(r,0,f_{1})\asymp(\log r)^{2}$ when $\lambda_{1}=0$ or $n(r,0,f_{1})\asymp r^{\lambda_{1}}$ when $0<\lambda_{1}\leq n$. Then number of zeros of $f_{2}$ has similar asymptotic behaviors. By slightly modifying the proof of Theorem 1.4, $f_{1}$ or $f_{2}$ can be chosen to have only finitely many zeros in the plane, as pointed out in Subsection 3.1.
As in [19, Theorem 3], we can describe the asymptotic behaviors of the Bank–Laine function $E$ in Theorem 1.3 and also the corresponding coefficient function $A$ of the second order differential equation (1.1) in the case $n=1$. We have the following
Theorem 1.5.
When $n=1$, the functions $A$ and $E$ in Theorem 1.4 can be chosen so that
$\begin{split}\frac{1}{2}\log|A(re^{i\theta})|\sim\log\frac{1}{|E(re^{i\theta})|}\sim\frac{1}{l}r\cos(\theta)\quad\text{for}\quad|\theta|\leq(1-\varepsilon)\frac{\pi}{2},\end{split}$
where $l$ is an integer such that $l=1$ when $\lambda_{1}\leq\lambda_{2}<1$ and $l=3$ when $\lambda_{1}<\lambda_{2}=1$ and $l=4$ when $\lambda_{1}=\lambda_{2}=1$, while
$\begin{split}\log|E(-re^{i\theta})|\sim\frac{1}{l}r\cos(\theta)\quad\text{for}\quad|\theta|\leq(1-\varepsilon)\frac{\pi}{2}\end{split}$
and
$\begin{split}|A(-re^{i\theta})|\sim\frac{1}{4l^{2}}\quad\text{for}\quad|\theta|\leq(1-\varepsilon)\frac{\pi}{2}\end{split}$
uniformly as $r\to\infty$, for any $\varepsilon\in(0,1)$.
The proof for Theorem 1.5 is only a slight modification of that in [19, Theorem 3] for the case $\rho=\sigma=1$ and thus will be omitted. The proof is based on the fact that the Schwarzian derivative $S(F)$ can be factored as $S(F)=B(F/F^{\prime})/2$, where
$\begin{split}B(E):=-2\frac{E^{\prime\prime}}{E}+\left(\frac{E^{\prime}}{E}\right)^{2}-\frac{1}{E^{2}}.\end{split}$
See Bank and Laine [2, 3]. The general solution of the differential equation $B(E)=4A$, that is, of the equation
(1.8) $\begin{split}4A=-2\frac{E^{\prime\prime}}{E}+\left(\frac{E^{\prime}}{E}\right)^{2}-\frac{1}{E^{2}}\end{split}$
is a product of two normalized solutions of (1.1). The results in [19, Theorem 3] shows that the functions $A$ and $E$ constructed in Theorem 1.3 are actually of completely regular growth (in the sense of Levin and Pfluger). This is obvious for $E$ and $A$ when $\lambda_{1}\leq\lambda_{2}<n$. In fact, by the classical Hadamard factorization theorem, Theorem 1.5 together with its extension to the general case in Subsection 3.4 implies that $E(z)=H(z)e^{p(z)}$, where $H$ is the canonical product formed by the zeros of $E$ and $p$ is a polynomial of degree $n$ of the form $p(z)=z^{n}+\cdots$.
While the Bank–Laine functions $E=f_{1}f_{2}$ constructed in previous three theorems all satisfy $\rho(E)<\infty$, by the linear independence of $f_{1}$ and $f_{2}$, the product $E=f_{1}(cf_{1}+f_{2})$ with a nonzero constant $c$ is also a Bank–Laine function and satisfies $\lambda(E)=\infty$. However, we shall prove the following
Theorem 1.6.
Let $\rho\in(1/2,\infty)$ and $\lambda\in[0,\infty)$. Then there exists an entire function $A$ of order $\rho$ such that the associated Bank–Laine function $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})=\lambda$ and $\log n(r,0,f_{2})\asymp r^{\rho}$. Moreover, the Bank–Laine function $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other solution $f_{3}$ linearly independent from $f_{1}$.
To prove Theorem 1.6, we not only glue the functions $g_{m,n}$ as in the proof of Theorem 1.2 but also glue the two functions $\hat{g}_{m,n}=(g_{m,n}+1)/2$ and $\hat{g}_{m,n}=g_{m,n}$ along horizontal lines for certain pair $(m,2n)$ of two nonnegative integers $m$ and $n$. If we fix $\check{g}_{m_{1},n_{1}}=g_{m_{1},n_{1}}$ for one pair $(m_{1},2n_{1})$ of two nonnegative integers $m_{1}$ and $n_{1}$ in the lower half-plane and $\hat{g}_{m_{2},n_{2}}=(g_{m_{2},n_{2}}+1)/2$ for another pair $(m_{2},2n_{2})$ of two nonnegative integers $m_{2}$ and $n_{2}$ in the upper half-plane, then we may follow the same process as in the proof of Theorem 1.3 to prove: for a dense set of orders $\rho\in(1,\infty)$, there exists an entire function $A$ of order $\rho(A)=\rho$ such that the associated Bank–Laine function $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})<\infty$ and $\lambda(f_{2})=\infty$. Moreover, the Bank–Laine function $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other solution $f_{3}$ linearly independent from $f_{1}$.
By the proof of Theorem 1.6, the constructed entire function $A$ and the corresponding Bank–Laine function $E$ inherit all the properties of those in Theorem 1.2 in the lower half-plane in the case $\rho\in(1/2,1)$, including the asymptotic behaviors as in [19, Theorem 3]. In fact, by checking the proof for [19, Theorem 3] together with the definition $\hat{g}_{m,n}=(g_{m,n}+1)/2$, we easily find that these asymptotic behaviors also hold for $z=re^{i\theta}$ such that $|\theta|<(1-\varepsilon)\pi/(2\sigma)$ with $\sigma=\rho/(2\rho-1)$. However, in the region $\{z=re^{i\theta}:0\leq\theta\leq\pi/(2\rho)\}$ of the plane, the function $E$ has much complicated behavior. There must be an infinite sequence $(z_{n})$, $z_{n}=r_{n}e^{i\theta_{n}}$, $0\leq\theta_{n}\leq\pi/(2\rho)$ and $r_{n}\to\infty$, such that $\log|E(r_{n}e^{i\theta_{n}})|\geq cr_{n}^{\rho}$ for some positive constant $c$ and all large $r_{n}$.
Theorem 1.6 can be viewed as an extension of Theorem 1.2. Then a combination of the methods in the proof of Theorem 1.4 and Theorem 1.6 yields an extension of Theorem 1.4, namely the following
Theorem 1.7.
Let $n\in\mathbb{N}$ be a positive integer and $\lambda_{1},\lambda_{2}\in[0,n]$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$ and $\lambda_{3}\in[0,\infty)$. Then there exists an entire function $A$ of order $n$ such that the associated Bank–Laine function $E=f_{1}f_{2}$ for two normalized solutions $f_{1}$ and $f_{2}$ of (1.1) satisfies $\lambda(f_{1})=\lambda_{1}$ and $n(r,0,f_{2})\asymp r^{\lambda_{3}}$ or $\log n(r,0,f_{2})\asymp(\log r)^{2}$ when $\lambda_{2}=0$ or $\log n(r,0,f_{2})\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq n$. Moreover, the Bank–Laine function $E=f_{1}f_{3}$ never satisfies $\lambda(E)<\infty$ for any other solution $f_{3}$ linearly independent from $f_{1}$.
We note that the Bank–Laine functions constructed in Theorem 1.6 and Theorem 1.7 all have infinite order. However, in Theorem 1.7 where $\rho(A)$ is an integer, each of the three cases $\lambda(E)>\rho(A)$, $\lambda(E)=\rho(A)$ or $\lambda(E)<\rho(A)$ can occur. Together with Theorem 1.4, this shows that the Bank–Laine function $E$ associated with an entire coefficient $A$ of integer order indeed have value distribution properties not covered by Theorem 1.1.
A question concerning the entire function $A$ in the second order differential equation (1.1) was asked in [9, Question 4.5]: Is it possible to characterize the transcendental entire functions $A$ where (1.1) admits a non-trivial solution $f_{1}$ with $\lambda(f_{1})<\infty$, such that for every solution $f_{2}$ of (1.1) that is linearly independent with $f_{1}$ we have $\lambda(f_{2})=\infty$? The constructions in Theorem 1.6 and Theorem 1.7 together the comments above present such entire functions $A$ of finite order. With modifications of the methods there, we may construct more such coefficient functions $A$ with certain growth properties.
In the remainder of this paper, we provide the proof for Theorem 1.3, Theorem 1.4, Theorem 1.6 and Theorem 1.7 in Section 2-Section 5, respectively. In the following sections, we keep the permanent notation: $(f_{1},f_{2})$ is a normalized pair of solutions of (1.1), the quotient $F=f_{2}/f_{1}$ is a solution of $S(F)=2A$, and $E=f_{1}f_{2}=F/F^{\prime}$ is a solution of (1.8). The notations $\phi,\psi,\tau,F,G$ and so on in each section will represent different functions.
2. Proof of Theorem 1.3
For every pair $(m,2n)$ of two nonnegative integers $m$ and $n$, we consider the rational function $R_{m,n}(z)$ in (1.4). In this section, we fix two distinct pairs $(m_{1},2n_{1})$ and $(m_{2},2n_{2})$ and will sometimes omit them from notation. We shall assume that at least one of $m_{1}$ and $m_{2}$ is nonzero and at least one of $n_{1}$ and $n_{2}$ is nonzero. Note that, unlike in [4], $g_{m_{1},n_{1}}$ and $g_{m_{2},n_{2}}$ are locally univalent meromorphic functions. The process of gluing two different $g_{m,n}$s is almost the same as that in [4]: by restricting $g_{m_{1},n_{1}}$ to the lower half-plane $\check{H}$ and $g_{m_{2},n_{2}}$ to the upper half-plane $\hat{H}$, we then paste them together, using a quasiconformal surgery, producing a meromorphic function $F$. Then our Bank–Laine function will be $E=F/F^{\prime}$ and thus $A=B(E)/4$ as in (1.8). For, we will use some different notations from the ones used in [5].
By the properties of $g_{m,n}$ described in the introduction, we see that there exists an increasing diffeomorphism $\phi:\mathbb{R}\to\mathbb{R}$ such that
(2.1) $g_{m_{2},n_{2}}(x)=(g_{m_{1},n_{1}}\circ\phi)(x)$
for all $x\in\mathbb{R}$. Denote $M=m_{2}+2n_{2}+1$ and $N=m_{1}+2n_{1}+1$ for simplicity. Let
$\begin{split}K=\frac{M}{N}=\frac{m_{2}+2n_{2}+1}{m_{1}+2n_{1}+1}.\end{split}$
Though $g_{m,n}$ in (1.5) has some poles, we may use the arguments as in [4] to show
Lemma 2.1.
The asymptotic behaviors of the diffeomorphism $\phi$ satisfy:
(2.2) $\phi(x)$ $=$ $x+O(e^{-x/2}),\ \ \ \ \ \ \ \ \ \ \ \ \ \phi^{\prime}(x)=1+O(e^{-x/2}),\quad\quad x\to\infty,$
(2.3) $\phi(x)$ $=$ $Kx+c+O(e^{-\delta|x|/2}),\ \ \ \ \phi^{\prime}(x)=K+O(e^{-\delta|x|/2}),\quad x\to-\infty,$
with
$\begin{split}c=\frac{1}{N}\log\left[\frac{\binom{m_{1}+2n_{1}}{m_{1}}}{\binom{m_{2}+2n_{2}}{m_{2}}}\frac{N!}{M!}\right]\quad\text{and}\quad\delta=\frac{1}{2}\min\{1,K\}.\end{split}$
Proof.
In order to prove (2.2), we note that
$\begin{split}\log g_{m_{2},n_{2}}(x)=e^{x}+O(x)=e^{x}(1+O(xe^{-x})),\quad x\to\infty.\end{split}$
The equation $g_{m_{2},n_{2}}(x)=(g_{m_{1},n_{1}}\circ\phi)(x)$ easily implies that $2x/3\leq\phi(x)\leq 2x$ for large $x$. Thus we also have
$\begin{split}\log g_{m_{1},n_{1}}(\phi(x))&=e^{\phi(x)}(1+O(\phi(x)e^{-\phi(x)}))=e^{\phi(x)}(1+O(\phi(x)e^{-2x/3})),\quad x\to\infty.\end{split}$
Combining the last two equations we obtain
$\begin{split}e^{\phi(x)-x}=1+O(xe^{-2x/3}),\quad x\to\infty,\end{split}$
from which the first statement in (2.2) easily follows. For the second statement in (2.2) we use
(2.4) $\begin{split}\phi^{\prime}=\frac{g^{\prime}_{m_{2},n_{2}}}{g_{m_{2},n_{2}}}\frac{g_{m_{1},n_{1}}\circ\phi}{g^{\prime}_{m_{1},n_{1}}\circ\phi}=e^{Mx-N\phi(x)}\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}\frac{\sum_{i=0}^{m_{1}}A_{i}e^{i\phi(x)}}{\sum_{i=0}^{m_{2}}A_{i}e^{ix}}\frac{\sum_{j=0}^{2n_{1}}B_{j}e^{j\phi(x)}}{\sum_{j=0}^{2n_{2}}B_{j}e^{jx}},\end{split}$
so that, together with the first statement in (2.2),
(2.5) $\begin{split}\phi^{\prime}&=e^{Mx-N\phi(x)+(m_{1}+2n_{1})\phi(x)-(m_{2}+2n_{2})x}\frac{1+O(e^{-\phi(x)})}{1+O(e^{-x})}\frac{1+O(e^{-\phi(x)})}{1+O(e^{-x})}\\ &=e^{O(e^{-x/2})}(1+O(e^{-\phi(x)})+O(e^{-x}))=1+O(e^{-x/2}),\quad x\to\infty.\end{split}$
In order to prove (2.3) we note from (1.5) and (1.6) that
$\begin{split}h^{\prime}_{m,n}(x)=\frac{1}{\binom{m+2n}{m}(m+2n)!}\frac{e^{x}x^{m+2n}}{Q_{m}(x)^{2}}.\end{split}$
Since $h_{m,n}(x)$ is analytic on $[0,\infty)$, then by taking the derivatives of $h^{\prime}_{m,n}(x)$ successively we find that $h^{\prime}_{m,n}(0)=\cdots=h^{(m+2n)}_{m,n}(0)=0$. Thus by the formula for the error term of a Taylor series we have
$\begin{split}h_{m,n}(x)=R_{m,n}(x)e^{x}=1+\frac{x^{m+2n+1}}{\binom{m+2n}{m}(m+2n+1)!}+O(x^{m+2n+2}),\quad x\to 0.\end{split}$
Hence
$\begin{split}g_{m,n}(x)&=1+\frac{e^{(m+2n+1)x}}{\binom{m+2n}{m}(m+2n+1)!}+O(e^{(m+2n+2)x})\\ &=1+\frac{e^{(m+2n+1)x}}{\binom{m+2n}{m}(m+2n+1)!}(1+O(e^{x})),\quad x\to-\infty.\end{split}$
The equation $g_{m_{2},n_{2}}(x)=g_{m_{1},n_{1}}(\phi(x))$ now yields
$\begin{split}\frac{\binom{m_{2}+2n_{2}}{m_{2}}}{\binom{m_{1}+2n_{1}}{m_{1}}}\frac{M!}{N!}e^{N\phi(x)-Mx}=1+O(e^{x})+O(e^{\phi(x)}),\quad x\to-\infty\end{split}$
and hence
$\begin{split}\phi(x)=\frac{M}{N}x+\frac{1}{N}\log\left[\frac{\binom{m_{2}+2n_{2}}{m_{2}}}{\binom{m_{1}+2n_{1}}{m_{1}}}\frac{M!}{N!}\right]+O(e^{x})+O(e^{\phi(x)}),\quad x\to-\infty,\end{split}$
which gives the first statement in (2.3). For the second statement in (2.3), recall that $A_{m}B_{2n}=\frac{m!(2n)!}{[(m+2n)!]^{2}}=\frac{1}{\binom{m+2n}{m}(m+2n)!}$. Similarly as for the estimates in (2.5), we use (2.4) together with the first statement in (2.3) to obtain
$\begin{split}\phi^{\prime}(x)&=\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}e^{Mx-N\phi(x)}\frac{1+O(e^{\phi(x)})}{1+O(e^{x})}\frac{1+O(e^{\phi(x)})}{1+O(e^{x})}\\ &=\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}e^{Mx-N(Kx+c+O(e^{-\delta|x|/2}))}(1+O(e^{\phi(x)})+O(e^{x}))\\ &=\frac{A_{m_{2}}B_{2n_{2}}}{A_{m_{1}}B_{2n_{1}}}e^{-Nc+O(e^{-\delta|x|/2})}(1+O(e^{\phi(x)})+O(e^{x}))=K+O(e^{-\delta|x|/2}).\end{split}$
Now we begin to define a quasiregular function in the same way as in [5]. Though the functions $g_{m_{1},n_{1}}$ and $g_{m_{2},n_{2}}$ we use here may have some poles, the process of definition here is in many parts the same as that in [5]. For the convenience of calculating the number of zeros of the desired Bank–Laine functions in the final part of the proof and also of the applications of the arguments in the proof of Theorem 1.4, Theorem 1.6 and Theorem 1.7, below we shall keep as many as details.
Let $D=\mathbb{C}\setminus\mathbb{R}_{\leq 0}$, and $p:D\to\mathbb{C}$, $p(z)=z^{\mu}$, the principal branch of the power. Here $\mu$ is a complex number such that
(2.6) $\begin{split}\mu=\text{Re}(\mu)+i\text{Im}(\mu)=\frac{2\pi}{4\pi^{2}+\log^{2}K}(2\pi-i\log K).\end{split}$
Note that $K\not=1$ under assumption. As shown in [5], $p$ maps $D$ onto the complement $G$ of a logarithmic spiral $\Gamma$, with
(2.7) $\begin{split}p(x+i0)=p(Kx-i0),\quad x<0.\end{split}$
The inverse map $h=p^{-1}$ is a conformal homeomorphism $h:G\to D$. Let $\Gamma^{\prime}=p(\mathbb{R}_{\geq 0})$. The two logarithmic spirals $\Gamma$ and $\Gamma^{\prime}$ divide the plane into two parts $\hat{G}$ and $\check{G}$, which are images under $p$ of the upper and lower half-planes, respectively.
The function $V$ defined by
$V(z)=\left\{\begin{array}[]{ll}(g_{m_{2},n_{2}}\circ h)(z),&\quad z\in\hat{G},\\ (g_{m_{1},n_{1}}\circ h)(z),&\quad z\in\check{G}\end{array}\right.$
is meromorphic in $\hat{G}\cup\check{G}$ and has a jump discontinuity on $\Gamma$ and $\Gamma^{\prime}$. To remove this discontinuity, we consider the strip $\Pi=\{z:|\text{Im}\ z|<1\}$ and define a quasiconformal homeomorphism $\psi:\mathbb{C}\to\mathbb{C}$ to be the identity outside of $\Pi$ and for $y=\text{Im}\ z\in(-1,1)$,
(2.8) $\psi(x+iy)=\left\{\begin{array}[]{ll}\phi(x)+|y|(x-\phi(x))+iy,&\quad x\geq 0,\\ \phi(x/K)+|y|(x-\phi(x/K)+iy),&\quad x<0.\end{array}\right.$
Then $\psi$ commutes with the complex conjugation, and satisfies
(2.9) $\begin{split}\psi(x)=\phi(x),\quad x>0,\quad\text{and}\quad\psi(Kx)=\phi(x),\quad x<0.\end{split}$
The Jacobian matrix $D_{\psi}$ of $\psi$ is given for $x>0$ and $10<|y|<$ 1 by
$D_{\psi}(x+iy)=\left(\begin{array}[]{cc}\phi^{\prime}(x)+|y|(1-\phi(x))&\pm(x-\phi(x))\\ 0&1\\ \end{array}\right),$
and we see using (2.2) that
$D_{\psi}(x+iy)=\left(\begin{array}[]{cc}1&0\\ 0&1\\ \end{array}\right),\quad 0<|y|<1,\quad x\to\infty.$
Similarly, using (2.3) we find that
$D_{\psi}(x+iy)=\left(\begin{array}[]{cc}1&\mp c\\ 0&1\\ \end{array}\right),\quad 0<|y|<1,\quad x\to-\infty.$
We conclude that $\psi$ is quasiconformal in the plane.
Now we modify $V$ to obtain a continuous function. Define $U:\mathbb{C}\to\mathbb{C}$ by
(2.10) $U(z)=\left\{\begin{array}[]{ll}(g_{m_{2},n_{2}}\circ h)(z),&\quad z\in\hat{G}\cup\Gamma\cup\Gamma^{\prime}\cup\{0\},\\ (g_{m_{1},n_{1}}\circ\psi\circ h)(z),&\quad z\in\check{G}.\end{array}\right.$
It follows from (2.7) and (2.8) that $U$ is continuous on $\Gamma$ and $\Gamma^{\prime}$ and quasiregular in the plane. The existence theorem for solutions of the Beltrami equation [14, § V.1] shows that there exists a quasiconformal homeomorphism $\tau:\mathbb{C}\to\mathbb{C}$ with the same Beltrami coefficient as $U$. The function $F=U\circ\tau^{-1}$ is then meromorphic. Moreover, as shown in [4], the Beltrami coefficient of $U$ (and hence of $\tau$) satisfies the hypotheses of the Techmüler–Wittich–Beliskii theorem [14, § V.6] (see also [17]). This theorem shows that $\tau$ is conformal at $\infty$ and may thus be normalized to satisfy
(2.11) $\begin{split}\tau(z)\sim z,\quad z\to\infty.\end{split}$
We set $E=F/F^{\prime}$. As $F^{\prime}(z)\not=0$ for all $z\in\mathbb{C}$ by construction, $E$ is entire. As all poles and zeros of $F$ are simple, all residues of $F/F^{\prime}$ are equal to $1$ or $-1$, so $E^{\prime}(z)=1$ or $E^{\prime}(z)=-1$ at every zero of $E$, which implies the Bank–Laine property. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A$.
We shall use a slightly different method from that in [4] to show that $E$, as well as the entire function $A$, has order $\rho=1/\text{Re}(\mu)$. Below we shall first estimate the numbers of poles and zeros of $F$ respectively. For two positive quantities $S_{1}(r)$ and $S_{2}(r)$, below we use the notation $S_{1}(r)\asymp S_{2}(r)$, provided that there are two positive constants $c_{1}$ and $c_{2}$ such that $c_{2}S_{2}(r)\leq S_{1}(r)\leq c_{1}S_{2}(r)$ for all large $r$.
For the meromorphic function $g_{m_{2},n_{2}}$, if $m_{2}\not=0$, then the number of poles of $g_{m_{2},n_{2}}$ satisfies $n(r,\infty,g_{m_{2},n_{2}})\sim m_{2}r/\pi$ as $r\to\infty$. Thus, by the periodicity of $g_{m_{2},n_{2}}$, the number of poles of $g_{m_{2},n_{2}}$ in the upper half-pane $\hat{H}$ satisfies $\hat{n}(r,\infty,g_{m_{2},n_{2}})\sim m_{2}r/(2\pi)$ as $r\to\infty$. We note that the image of the circle $\{z:|z|=r\}$ under $h(z)=z^{1/\mu}$ is the part of the logarithmic spiral which connects two points on the negative real axis and intersects the positive real axis at $r^{1/\text{Re}(\mu)}$. By the definitions of $p$ and $h$ we may write $z=re^{i\theta}$ on the circle such that $z_{0}=re^{i\theta_{0}}$ is a point on $\Gamma^{\prime}$ and
$\begin{split}h(z)=\exp\left(\frac{1}{\mu}\log z\right)=\exp\left(\frac{(\text{Im}(\mu))\theta+(\text{Re}(\mu))\log r}{|\mu|^{2}}+i\frac{(\text{Re}(\mu))\theta-(\text{Im}(\mu))\log r}{|\mu|^{2}}\right).\end{split}$
When $K>1$, noting that $\text{Re}(\mu)>0$ and $\text{Im}(\mu)<0$, we may write $-(\theta_{0}+\pi)\leq\theta\leq-(\theta_{0}-\pi)$, where $\theta_{0}$ satisfies $(\text{Re}(\mu))\theta_{0}=(\text{Im}(\mu))\log r$. A simple computation shows that
(2.12) $\begin{split}e^{(\text{Im}(\mu))\pi}r^{1/\text{Re}(\mu)}\leq|h(re^{i\theta})|\leq e^{-(\text{Im}(\mu))\pi}r^{1/\text{Re}(\mu)}.\end{split}$
Therefore, when $r$ is large, the number of poles of $g_{m_{2},n_{2}}\circ h$ in $\hat{G}$ satisfies
$\begin{split}c_{2}r^{1/\text{Re}(\mu)}(\log r)\leq\hat{n}(r,\infty,g_{m_{2},n_{2}}\circ h)\leq c_{1}r^{1/\text{Re}(\mu)}(\log r)\end{split}$
for two positive constants $c_{1}$ and $c_{2}$ both dependent on $K$. Similarly, if $m_{1}\not=0$, the number of poles of $g_{m_{1},n_{1}}\circ h$ in $\check{G}$ satisfies
(2.13) $\begin{split}c_{4}r^{1/\text{Re}(\mu)}(\log r)\leq\check{n}(r,\infty,g_{m_{1},n_{1}}\circ h)\leq c_{3}r^{1/\text{Re}(\mu)}(\log r)\end{split}$
for two positive constants $c_{3}$ and $c_{4}$ both dependent on $K$. By the definition of $\tau(x+iy)$ in (2.8) and the asymptotic behaviors for $\phi$ in (2.2) and (2.3), we see that the number of poles of $g_{m_{1},n_{1}}\circ\psi\circ h$ in $\check{G}$ satisfies
(2.14) $\begin{split}c_{6}r^{1/\text{Re}(\mu)}(\log r)\leq\check{n}(r,\infty,g_{m_{1},n_{1}}\circ\psi\circ h)\leq c_{5}r^{1/\text{Re}(\mu)}(\log r)\end{split}$
for two positive constants $c_{5}$ and $c_{6}$ both dependent on $K$. When $K<1$, we can also obtain two similar estimates as in (2.13) and (2.14). By summarizing the above estimates together with the relation in (2.11), we conclude that $n(r,\infty,F)\asymp r^{1/\text{Re}(\mu)}$. It follows that $N(r,\infty,F)\asymp r^{1/\text{Re}(\mu)}$. Similarly, when at least one of $n_{1}$ and $n_{2}$ is nonzero, for the zeros of $F$ we also obtain $n(r,0,F)\asymp r^{1/\text{Re}(\mu)}$ and thus $N(r,0,F)\asymp r^{1/\text{Re}(\mu)}$. Since $E=F/F^{\prime}$ and since the poles and zeros of $F$ are simple, we have
(2.15) $\begin{split}N\left(r,0,E\right)\asymp r^{1/\text{Re}(\mu)}(\log r).\end{split}$
Next we estimate $m(r,1/E)$. We note that there is a set $\Omega_{1}$ of finite linear measure such that, for all $|z|\not\in\Omega_{1}$, the functions $P_{n}(e^{z})$ and $Q_{m}(e^{z})$ both satisfy
$\left\{\begin{array}[]{ll}\exp(-(2n_{i}+\varepsilon)|z|)\leq\left|P_{n_{i}}(e^{z})\right|\leq\exp((2n_{i}+\varepsilon)|z|),&\\ \exp(-(m_{i}+\varepsilon)|z|)\leq\left|Q_{m_{i}}(e^{z})\right|\leq\exp((m_{i}+\varepsilon)|z|),&\end{array}\right.$
where $i=1,2$ and $\varepsilon>0$ is a small constant. This implies that
$\begin{split}|g_{m_{i},n_{i}}(z)|\leq\exp((m_{i}+\varepsilon)|z|+(2n_{i}+\varepsilon)|z|)\left|\exp(e^{z})\right|\end{split}$
for $i=1,2$ and all $z\in\mathbb{C}$ such that $|z|\not\in\Omega_{1}$. Denote $L_{i}=m_{i}+2n_{i}$, $i=1,2$, for simplicity. Clearly, this implies that there is a set $\Omega_{2}$ of finite linear measure such that
(2.16) $\left\{\begin{array}[]{ll}|g_{m_{2},n_{2}}(z)|\leq\exp((L_{2}+2\varepsilon)|z|)\left|\exp\left(e^{z}\right)\right|\quad\text{for}\quad z\in H^{+}\quad\text{and}\quad|z|\not\in\Omega_{2},&\\ |g_{m_{1},n_{1}}(z)|\leq\exp((L_{1}+2\varepsilon)|z|)\left|\exp\left(e^{z}\right)\right|\quad\text{for}\quad z\in H^{-}\quad\text{and}\quad|z|\not\in\Omega_{2}.&\\ \end{array}\right.$
We deduce from (2.10), (2.12) and (2.16) that there is a set $\Omega_{3}$ of finite linear measure such that
$\begin{split}\log|g_{m_{2},n_{2}}\circ h|\leq c(L_{2}+2\varepsilon)|z|^{1/\text{Re}(\mu)}+\exp(c|z|^{1/\text{Re}(\mu)})\quad\text{for}\quad z\in G^{+}\quad\text{and}\quad|z|\not\in\Omega_{3},\end{split}$
where $c$ is a suitably chosen positive constant and, together with the asymptotic behaviors for $\phi$ in (2.2) and (2.3) and also the definition $\psi(x+iy)$ in (2.9), that
$\begin{split}\log|g_{m_{1},n_{1}}\circ\psi\circ h|\leq c(L_{1}+2\varepsilon)|z|^{1/\text{Re}(\mu)}+\exp(c|z|^{1/\text{Re}(\mu)})\quad\text{for}\quad z\in G^{-}\quad\text{and}\quad|z|\not\in\Omega_{3}.\end{split}$
Hence (2.11) together with the above two estimates yields that
(2.17) $\begin{split}\log|F(z)|\leq\exp((c+o(1))|z|^{1/\text{Re}(\mu)})\end{split}$
as $|z|\to\infty$ outside a set $\Omega_{3}$ of finite linear measure. Together with the relation in (2.17), the lemma on the logarithmic derivative [10, Section 2.2] now implies that $E=F/F^{\prime}$ satisfies
(2.18) $\begin{split}m\left(r,\frac{1}{E}\right)=O(r^{1/\text{Re}(\mu)}),\end{split}$
outside a set of finite linear measure. Thus, by (2.15) and the first main theorem of Nevanlinna theory [10, Section 1.3], the characteristic function $T(r,E)$ of $E$ satisfies
$\begin{split}T\left(r,E\right)\asymp r^{1/\text{Re}(\mu)}(\log r).\end{split}$
Therefore, $E=f_{1}f_{2}$ is a Bank–Laine function such that $\lambda(E)=\rho(E)=1/\text{Re}(\mu)=\rho$, as desired. Moreover, the two normalized solutions $f_{1}$ and $f_{2}$ of the second order linear differential equation (1.1) both satisfy $n(r,0,f_{i})\asymp r^{1/\text{Re}(\mu)}$, $i=1,2$.
Finally, we determine the order of the entire function $A$. Since $E$ is of finite order, it follows by the relation (1.8) together with (2.18) and the lemma on the logarithmic derivative that
$\begin{split}m(r,A)=m\left(r,\frac{1}{E}\right)+O(\log r)=O(r^{1/\text{Re}(\mu)}).\end{split}$
Thus the first main theorem implies that $\rho(A)\leq 1/\text{Re}(\mu)$. On the other hand, we note that the function $U$ defined in (2.10) is meromorphic in $G^{+}$ and, in particular, is analytic on $\Gamma$ and $\Gamma^{\prime}$ respectively. By taking the derivative of $F$ in $\tau(\hat{G})$, we get
(2.19) $\begin{split}\frac{F^{\prime}}{F}=\left(\frac{g^{\prime}_{m_{2},n_{2}}}{g_{m_{2},n_{2}}}\circ h\circ\tau^{-1}\right)(h^{\prime}\circ\tau^{-1})(\tau^{-1})^{\prime}.\end{split}$
We may choose a curve in $\tau(\hat{G})$, say $\Gamma^{\prime\prime}$, so that $\tau^{-1}(\Gamma^{\prime\prime})$ is close to the curve $\Gamma^{\prime}$. For a point $z\in\tau^{-1}(\Gamma^{\prime\prime})$, we have the Cauchy formula
[…]
with a circle $C_{z}$ centered at $z$. Choosing the radius $\beta(z)$ of this circle to be finite, say $\beta(z)=1$, we have $|\tau^{\prime}(z)|\leq 2|z|$ when $|z|$ is large. Since $E=F/F^{\prime}$, then by the definition of $g_{m,n}$ in (1.5) as well as the relations in (2.11) and in (2.19), we see that $1/E$ satisfies
(2.20) $\begin{split}\frac{1}{|E(z)|}\geq\left(d_{1}e^{\frac{1}{2}|z|^{1/\text{Re}(\mu)}}\right)\left(\frac{d_{2}|z|^{1/\text{Re}(\mu)-1}}{|\mu|}\right)\frac{1}{2|z|}\end{split}$
for two suitably positive constants $d_{1},d_{2}$ and all $z\in\Gamma^{\prime\prime}$ such that $|z|$ is large. Further, since $E$ is of finite order, by an estimate in [7, Corollary 2] we may suppose that, for a small $\varepsilon>0$, there is an infinite sequence $(z_{n})$ such that $z_{n}\in\Gamma^{\prime\prime}$ and $|E^{\prime}(z_{n})/E(z_{n})|\leq|z_{n}|^{1/\text{Re}(\mu)-1+\varepsilon}$ and $|E^{\prime\prime}(z_{n})/E(z_{n})|\leq|z_{n}|^{2(1/\text{Re}(\mu)-1+\varepsilon)}$ for all $z_{n}$ such that $|z_{n}|$ is large. Together with these two estimates as well as the inequality in (2.20), the relation in (1.8) implies that $\log|A(z_{n})|\geq d|z_{n}|^{1/\text{Re}(\mu)}$ for some positive constant $d$ and all $z_{n}\in\Gamma^{\prime\prime}$ such that $|z_{n}|$ is large, implying that $\rho(A)\geq 1/\text{Re}(\mu)$. Hence $\rho(A)=1/\text{Re}(\mu)$. This completes the proof of Theorem 1.3.
3. Proof of Theorem 1.4
3.1. Preliminary lemmas
We first prove a technical lemma which will be used to control the number of zeros of the desired Bank–Laine functions. In this subsection, we restrict to consider the case $\lambda_{1}\leq\lambda_{2}<1$. We shall choose two sequences $(m_{k})$ and $(n_{k})$ associated with $\lambda_{1}$ and $\lambda_{2}$, respectively.
Lemma 3.1.
Let $\lambda\in[0,1)$. Then there is a sequence $(m_{k})$ of integers 0 and 1 only, with $m_{k}=0$ for $k=1,2,3$, such that the function $h:[0,\infty)\to[0,\infty)$ which satisfies $h(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ and has slopes $m_{k}$ there satisfies
(3.1) $\begin{split}h(x)=(\log x)^{2}+O(x^{-1}\log x)+O(x^{-1})+O(1),\quad x\to\infty\end{split}$
when $\lambda=0$, or
(3.2) $\begin{split}h(x)=x^{\lambda}+O(1),\quad x\to\infty\end{split}$
when $10<\lambda<$ 1. Moreover, when $\lambda=0$, the subsequence $(m_{k_{i}})$ such that $m_{k_{i}}=1$ for all $k_{i}$ satisfies
(3.3) $\begin{split}2\pi k_{i}=\left[1+O\left(1/k_{i}\right)\right]e^{(2\pi i)^{1/2}},\quad i\to\infty\end{split}$
and, when $10<\lambda<$ 1, the subsequence $(m_{k_{i}})$ such that $m_{k_{i}}=1$ for all $k_{i}$ satisfies
(3.4) $\begin{split}2\pi k_{i}=(2\pi i)^{1/\lambda}+O(1),\quad i\to\infty.\end{split}$
Proof.
The statement in [19, Lemma 2.2] contains an error, so here we present a complete proof for the lemma. Moreover, we will construct the sequence $(m_{k})$ including infinitely many elements $1$ for the case $\lambda=0$.
When $10\leq\lambda<$ 1, we set $h(0)=0$ and choose $m_{1}=m_{2}=m_{3}=0$. We note that
$\begin{split}(2\pi k)^{\lambda}-(2\pi(k-1))^{\lambda}\leq 2\pi\quad\text{for}\quad k\geq 1.\end{split}$
Suppose that $k\geq 3$, and $h(2\pi k_{i-1})$ is already defined. Then we define
(3.5) $\begin{split}h(2\pi k_{i}):=2\pi i,\end{split}$
where $k_{i}$ is the largest positive integer such that
(3.6) $\begin{split}[\log(2\pi(k_{i}-1))]^{2}<2\pi i\leq[\log(2\pi k_{i})]^{2}\end{split}$
when $\lambda=0$, and the largest positive integer such that
(3.7) $\begin{split}(2\pi(k_{i}-1))^{\lambda}<2\pi i\leq(2\pi k_{i})^{\lambda}\end{split}$
when $10<\lambda<$ 1. Then we interpolate $h$ linearly between $2\pi(k_{i}-1)$ and $2\pi k_{i}$ in the way that the slopes are equal to $m_{k_{i}}=1$ in the intervals $[2\pi(k_{i}-1),2\pi k_{i}]$ and the slopes are equal to $m_{k}=0$ otherwise. In particular, when $m_{k_{i}}=1$, since $2\pi(k_{i}-1)<x<2\pi k_{i}$ we have $x=2\pi k_{i}+O(1)$. Using the relation $\log(1+x)\to x$ as $x\to 0$, the assertion on $h$ in (3.1) and (3.2) follows. Moreover, the two estimates in (3.3) and (3.4) follow from (3.6) and (3.7), respectively.
We can of course choose the sequence $(m_{k})$ in Lemma 3.1 to be such that only finitely many of the elements $m_{k}$ are equal to 1. In this case, $f_{1}$ or $f_{2}$ in the constructed Bank–Laine function $E=f_{1}f_{2}$ has only finitely many zeros.
Let $\lambda_{1},\lambda_{2}\in[0,1)$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$. By Lemma 3.1, for the number $\lambda_{1}\in[0,1)$, there is a sequence $(m_{k})$ of 00 and $1$ only and a function $h_{1}:[0,\infty)\to[0,\infty)$ which satisfies $h_{1}(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ and has slopes $m_{k}$ there satisfies
$\begin{split}h_{1}(x)=(\log x)^{2}+O(x^{-1}\log x)+O(x^{-1})+O(1),\quad x\to\infty\end{split}$
when $\lambda_{1}=0$, and satisfies
$\begin{split}h_{1}(x)=x^{\lambda_{1}}+O(1),\quad x\to\infty\end{split}$
when $10<\lambda_{1}<$ 1. Moreover, when $\lambda_{1}=0$, the subsequence $(m_{k_{i}})$ such that $m_{k_{i}}=1$ for all $k_{i}$ satisfies
$\begin{split}2\pi k_{i}=\left[1+O\left(1/k_{i}\right)\right]e^{(2\pi i)^{1/2}},\quad i\to\infty\end{split}$
and, when $10<\lambda_{1}<$ 1, the subsequence $(m_{k_{i}})$ such that $m_{k_{i}}=1$ for all $k_{i}$ satisfies
$\begin{split}2\pi k_{i}=(2\pi i)^{1/\lambda_{1}}+O(1),\quad i\to\infty.\end{split}$
For the number $\lambda_{2}\in[0,1)$, there is a sequence $(n_{k})$ of 00 and $1$ only and a function $h_{2}:[0,\infty)\to[0,\infty)$ which satisfies $h_{2}(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ and has slopes $2n_{k}$ there satisfies
$\begin{split}h_{2}(x)=2(\log x)^{2}+O(x^{-1}\log x)+O(x^{-1})+O(1),\quad x\to\infty\end{split}$
when $\lambda_{2}=0$, and satisfies
$\begin{split}h_{2}(x)=2x^{\lambda_{2}}+O(1),\quad x\to\infty\end{split}$
when $10<\lambda_{2}<$ 1. Moreover, when $\lambda_{2}=0$, the subsequence $(n_{k_{j}})$ such that $n_{k_{j}}=1$ for all $k_{j}$ satisfies
$\begin{split}2\pi k_{j}=\left[1+O\left(1/k_{j}\right)\right]e^{(2\pi j)^{1/2}},\quad j\to\infty\end{split}$
and, when $10<\lambda_{2}<$ 1, the subsequence $(n_{k_{j}})$ such that $n_{k_{j}}=1$ for all $k_{j}$ satisfies
$\begin{split}2\pi k_{j}=(2\pi j)^{1/\lambda_{2}}+O(1),\quad j\to\infty.\end{split}$
For each pair $(m_{k},2n_{k})$ of two nonnegative integers $m_{k}$ and $n_{k}$, we denote $N_{k}=m_{k}+2n_{k}+1$. Then, for the two sequences $(m_{k})$ and $(n_{k})$ chosen, a simple calculation shows that the function $H:[0,\infty)\to[0,\infty)$ which satisfies $H(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ and has slopes $N_{k}$ there satisfies
(3.8) $\begin{split}H(x)=x+\omega(x)\end{split}$
where $\omega:[0,\infty)\to[0,\infty)$ is a function satisfying $\omega(x)\sim 2(\log x)^{2}$ as $x\to\infty$ when $\lambda_{1}=\lambda_{2}=0$ or $\omega(x)\sim 2x^{\lambda_{2}}$ as $x\to\infty$ when $\lambda_{1}<\lambda_{2}$ or $\omega(x)\sim 3x^{\lambda_{2}}$ as $x\to\infty$ when $\lambda_{1}=\lambda_{2}>0$. It is easy to see that the distance between the curve determined by the function $y=H(x)$ and the curve determined by the function $y=x$ tends to infinity as $x\to\infty$. This makes it impossible to find a quasiconformal mapping as the one $Q$ in the plane to glue as in [19, Subsection 2.2].
We shall expand the intervals $[2\pi(k-1),2\pi k]$ suitably to obtain a function $g:[0,\infty)\to[0,\infty)$ which satisfies $g(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ such that
$\begin{split}H(g(x))=x.\end{split}$
Such a function $g$ exists and, by the estimates in (3.8), we see that $g(x)=x(1+o(1))$ as $x\to\infty$. To determine one such $g$ explicitly, we denote $\mathcal{N}_{k}=\sum_{j=1}^{k}N_{j}$ and modify the previous functions $h_{1}$, $h_{2}$ and $H$ in the intervals $[2\pi\mathcal{N}_{k-1},2\pi\mathcal{N}_{k}]$ to be linear there and has slops $m_{k}/N_{k}$, $2n_{k}/N_{k}$ and $1$, respectively. Then, for the constant $\lambda_{1}\in[0,1)$ and the sequence $(m_{k})$ before, there is a function $H_{1}:[0,\infty)\to[0,\infty)$ which satisfies $H_{1}(0)=0$ and which is linear in the intervals $[2\pi\mathcal{N}_{k-1},2\pi\mathcal{N}_{k}]$ and has slopes $m_{k}/N_{k}$ there satisfies
$\begin{split}H_{1}(x)=h_{1}(x)+O(1),\quad x\to\infty.\end{split}$
Moreover, when $\lambda_{1}=0$, the subsequence $(m_{k_{i}})$ such that $m_{k_{i}}=1$ for all $k_{i}$ satisfies
(3.9) $\begin{split}c_{2}e^{(2\pi i)^{1/2}}\leq 2\pi k_{i}\leq c_{1}e^{(2\pi i)^{1/2}}\end{split}$
for two positive constants $c_{1},c_{2}$ and, when $10<\lambda_{1}<$ 1, the subsequence $(m_{k_{i}})$ such that $m_{k_{i}}=1$ for all $k_{i}$ satisfies
(3.10) $\begin{split}c_{4}(2\pi i)^{1/\lambda_{1}}\leq 2\pi k_{i}\leq c_{3}(2\pi i)^{1/\lambda_{1}}\end{split}$
for two positive constants $c_{3},c_{4}$. Also, for the constant $\lambda_{2}\in[0,1)$ and the sequence $(n_{k})$ before, there is a function $H_{2}:[0,\infty)\to[0,\infty)$ which satisfies $H_{2}(0)=0$ and which is linear in the intervals $[2\pi\mathcal{N}_{k-1},2\pi\mathcal{N}_{k}]$ and has slopes $2n_{k}/N_{k}$ there satisfies
$\begin{split}H_{2}(x)=h_{2}(x)+O(1),\quad x\to\infty.\end{split}$
Moreover, when $\lambda_{2}=0$, the subsequence $(n_{k_{j}})$ such that $n_{k_{j}}=1$ for all $k_{j}$ satisfies
(3.11) $\begin{split}c_{6}e^{(2\pi j)^{1/2}}\leq 2\pi k_{j}\leq c_{5}e^{(2\pi j)^{1/2}}\end{split}$
for two positive constants $c_{5},c_{6}$ and, when $10<\lambda_{2}<$ 1, the subsequence $(n_{k_{j}})$ such that $n_{k_{j}}=1$ for all $k_{j}$ satisfies
(3.12) $\begin{split}c_{8}(2\pi j)^{1/\lambda_{2}}\leq 2\pi k_{j}\leq c_{7}(2\pi j)^{1/\lambda_{2}}\end{split}$
for two positive constants $c_{7},c_{8}$.
Let now $m,n,\mathfrak{m},\mathfrak{n}\in\mathbb{N}$ and $N=m+2n+1$ and $M=\mathfrak{m}+2\mathfrak{n}+1$. Let
$\begin{split}K=\frac{M}{N}=\frac{\mathfrak{m}+2\mathfrak{n}+1}{m+2n+1}.\end{split}$
Then we need to study the asymptotic behaviors of the increasing diffeomorphism $\phi:\mathbb{R}\to\mathbb{R}$ defined by
(3.13) $\begin{split}g_{\mathfrak{m},\mathfrak{n}}(x)=(g_{m,n}\circ\phi)(x),\end{split}$
which has appeared in (2.1). We will consider the functions $\phi$ for the case that $m=m_{k}$, $n=n_{k}$, $\mathfrak{m}=m_{k+1}$ and $\mathfrak{n}=n_{k+1}$ and also that $N=N_{k}$ and $M=N_{k+1}$. Note that $N_{k}\leq 4$ for all $k$ and thus $K_{k}$ is bounded for all $k$. Thus the asymptotic behaviors described in Lemma 2.1 apply to all $\phi_{k}$ such that $g_{m_{k+1},n_{k+1}}(x)=g_{m_{k},n_{k}}(\phi_{k}(x))$. In order to simplify the formulas, we will sometimes suppress the dependence of $\phi$ from $m,n$ and $\mathfrak{m},\mathfrak{n}$ from the notation.
For technical reasons, we will not work with $g_{m,n}$ defined in (1.5) directly. Recall that by (1.6) the function $g_{m,n}:\mathbb{R}\to(1,\infty)$ is an increasing homeomorphism. Thus there exists $s_{m,n}\in\mathbb{R}$ such that
(3.14) $\begin{split}g_{m,n}(s_{m,n})=2.\end{split}$
For each pair of $(m,2n)$ of two nonnegative integers $m$ and $n$, as in [5] we shall work with the function $g_{m,n}(z+s_{m,n})$. Since $N\leq 4$ in any case, we have the following
Lemma 3.2.
The constant $s_{m,n}$ in (3.14) satisfies $s_{m,n}=O(1)$ as $k\to\infty$.
Proof.
The proof for large $N$ has been given in [19, Lemma 2.4]. We write $s_{m,n}=\log N+r$. Then equation (1.5) yields
(3.15) $\begin{split}\log\left(\sum_{j=0}^{2n}B_{j}N^{j}e^{jr}\right)+\log\left(\sum_{i=0}^{m}A_{i}N^{i}e^{ir}\right)+Ne^{r}=\log 2.\end{split}$
The left hand side of (3.15) tends to 00 as $r\to-\infty$ and the left hand side of (3.15) tends to $\infty$ as $r\to\infty$. Therefore, $r=O(1)$.
3.2. Definition of a quasiregular map
Now we begin to define a quasiregular map by gluing the functions $g_{m,n}$ defined in (1.5) for two suitably chosen sequences $(m_{k})$ and $(n_{k})$. By modifying the functions $g_{m,n}$ suitably, we obtain closely related functions $u_{m,n}$ and $v_{m,n}$ and then glue restrictions of these maps to strips along horizontal lines to obtain quasiregular maps $U$ and $V$ which are defined in the right and left half-planes respectively. In particular, we have $U(iy)=V(iy)$ on the imaginary axis, resulting in a quasiregular map $G$ in the plane. The maps $U$, $V$ and $G$ will commute with the complex conjugation, so it will be enough to define them in the upper half-plane.
Instead of gluing the functions $g_{m,n}$ in (1.5) directly, for each pair $(m,2n)$ of nonnegative integers $m$ and $n$, we shall first modify the functions $g_{m,n}$ slightly and define
(3.16) $\begin{split}g^{+}_{m,n}(z)=g_{m,n}\left(\frac{x}{l}+i\frac{y}{N}\right),\end{split}$
where $l$ is a positive integer dependent on the sequence $(m_{k})$ and $(n_{k})$ we shall choose later, and
(3.17) $\begin{split}g^{-}_{m,n}(z)=g_{m,n}\left(\frac{x+iy}{N}\right).\end{split}$
Actually we shall restrict the functions in (3.16) to the right half-plane and the functions in (3.17) to the left half-plane. Thus we may write
(3.18) $\begin{split}\tilde{g}_{m,n}(z)=(g_{m,n}\circ\chi_{m,n})(z)\end{split}$
for a map $\chi_{m,n}:\mathbb{C}\to\mathbb{C}$ such that $\chi_{m,n}(x+iy)=x/l+iy/N$ in the right half-plane and $\chi_{m,n}(x+iy)=(x+iy)/N$ in the left half-plane respectively. Note that $\chi_{m,n}$ is an affine mapping when restricted to the right or left half-plane and also that $\chi_{m,n}$ is continuous on the imaginary axis.
Now, for the two numbers $\lambda_{1},\lambda_{2}\in[0,1]$ such that $\lambda_{1}\leq\lambda_{2}$, we choose the associated sequences $(m_{k})$ and $(n_{k})$ in the following way:
  • (I)
    if $\lambda_{1}\leq\lambda_{2}<1$, then we choose the sequence $(m_{k})$ and $(n_{k})$ constructed in Subsection 3.1 respectively;
  • (II)
    if $\lambda_{1}<\lambda_{2}=1$, then we choose the sequence $(m_{k})$ constructed in Subsection 3.1 and $(n_{k})$ in the way that $n_{k}=0$ for $k=1,2,3$ and $n_{k}=1$ for all $k\geq 4$ respectively;
  • (III)
    if $\lambda_{1}=\lambda_{2}=1$, then we choose the sequence $(m_{k})$ and $(n_{k})$ in the way that $m_{k}=0$ for $k=1,2,3$ and $m_{k}=1$ for all $k\geq 4$ and $n_{k}=0$ for $k=1,2,3$ and $n_{k}=1$ for all $k\geq 4$ respectively.
Corresponding to the two sequences $(m_{k})$ and $(n_{k})$ in each case, we shall choose the integer $l$ in $\tilde{g}_{m,n}$ in (3.18) as follows: $l=1$ in the case (I) and $l=3$ in the case (II) and $l=4$ in the case (III). Then the explicit expression for $\tilde{g}_{m,n}$ in (3.18) can be given using $g^{+}_{m,n}$ defined in (3.16) and $g^{-}_{m,n}$ defined in (3.17) as follows. On each strip
(3.19) $\begin{split}\hat{\mathbf{\Pi}}_{k}=\{x+iy:\quad y\geq 0,\quad 2\pi\mathcal{N}_{k-1}<y<2\pi\mathcal{N}_{k}\}\end{split}$
of the upper half-plane, we shall define
(3.20) $\begin{split}\tilde{g}_{m_{k},n_{k}}(z)=\left\{\begin{array}[]{ll}g_{m_{k},n_{k}}\left(\frac{x}{l}+i\frac{y-2\pi\mathcal{N}_{k-1}}{N_{k}}\right),&x\geq 0,\\ g_{m_{k},n_{k}}\left(\frac{x+iy-2i\pi\mathcal{N}_{k-1}}{N_{k}}\right),&x\leq 0.\end{array}\right.\end{split}$
Note that the two expressions in the definition of $\tilde{g}_{m_{k},n_{k}}$ in (3.20) coincide on the imaginary axis in each of three cases (I), (II) and (III).
Now we begin to glue the functions $\tilde{g}_{m,n}$ defined in (3.18) along the horizontal lines. We shall always assume that $(m_{k})$ and $(n_{k})$ are the two sequences chosen in either of the three cases. In either case, the choice of the integer $l$ guarantees that the resulting functions $U$, $V$ and $G$ are analytic in almost all parts of the plane.
We begin by constructing the map $U$. We shall use the explicit expressions for $\tilde{g}_{m_{k},n_{k}}$ in (3.20). Recall from equation (3.18) that $\tilde{g}_{m,n}=g_{m,n}\circ\chi_{m,n}$ for an affine mapping $\chi_{m,n}$ in the right half-plane. Instead of $\tilde{g}_{m,n}$, we consider the map
$\begin{split}u_{m,n}:\{z\in\mathbb{C}:\text{Re}\ z\geq 0\}\to\mathbb{C},\quad u_{m,n}(z)=g_{m,n}(\chi_{m,n}(z)+s_{m,n}).\end{split}$
Note that $u_{m,n}$ is increasing on the real line and maps $[0,\infty)$ onto $[2,\infty)$. For the two chosen sequences $(m_{k})$ and $(n_{k})$, we write $N_{k}=m_{k}+2n_{k}+1$. Put $U(z)=u_{m_{k},n_{k}}$ in the half-strip
(3.21) $\begin{split}\hat{\mathbf{\Pi}}_{k}^{+}=\{x+iy:x>0,\quad 2\pi\mathcal{N}_{k-1}<y<2\pi\mathcal{N}_{k}\}.\end{split}$
The function $U$ may be discontinuous along the horizontal lines $\mathcal{L}^{+}_{k-1}=\{z=x+iy:x>0,y=2\pi\mathcal{N}_{k-1}\}$ and $\mathcal{L}^{+}_{k}=\{z=x+iy:x>0,y=2\pi\mathcal{N}_{k}\}$ if $m_{k}=1$ or $n_{k}=1$. In order to obtain a continuous function we consider the function $\psi_{k}:[0,\infty)\to[0,\infty)$ defined by $u_{m_{k+1},n_{k+1}}(x)=u_{m_{k},n_{k}}(\psi_{k}(x))$. The function $\psi_{k}$ is closely related to the function $\phi$ considered in (3.13). In fact, denoting by $\phi_{k}$ the function $\phi$ corresponding to $(m_{k},2n_{k})$ and $(m_{k+1},2n_{k+1})$, we have
(3.22) $\begin{split}\psi_{k}(x)=l\phi_{k}\left(\frac{x}{l}+s_{m_{k+1},n_{k+1}}\right)-ls_{m_{k},n_{k}}.\end{split}$
We then define $U:\{z\in\mathbb{C}:\text{Re}\ z\geq 0\}\to\mathbb{C}$ by interpolating between $u_{m_{k+1},n_{k+1}}$ and $u_{m_{k},n_{k}}$ as follows: if $2\pi\mathcal{N}_{k-1}\leq y<2\pi\mathcal{N}_{k}$, say $y=2\pi\mathcal{N}_{k-1}+2\pi N_{k}t$ where $10\leq t<$ 1, then we put
$\begin{split}U(x+iy)=u_{m_{k},n_{k}}((1-t)x+t\psi_{k}(x)+iy)=u_{m_{k},n_{k}}(x+iy+t(\psi_{k}(x)-x)).\end{split}$
The function $U$ defined in this way is continuous along the horizontal lines $\mathcal{L}^{+}_{k}$ for any $k$ in the right-half plane.
We now define a function $V$ in the left half-plane. In order to do so, we define
$\begin{split}v_{m,n}:\{z\in\mathbb{C}:\text{Re}\ z<0\}\to\mathbb{C},\quad v_{m,n}(z)=g_{m,n}\left(\chi_{m,n}(z)+s_{m,n}\right).\end{split}$
Note that $v_{m,n}$ maps $(-\infty,0]$ monotonically onto $(1,2]$. For the two sequences $(m_{k})$ and $(n_{k})$ as before, this time we would like to define $V(z)=v_{m_{k},n_{k}}(z)$ in the half-strip
[…]
but again this function would be discontinuous along the horizontal lines $\mathcal{L}^{-}_{k-1}=\{z=x+iy:x<0,y=2\pi\mathcal{N}_{k-1}\}$ and $\mathcal{L}^{-}_{k}=\{z=x+iy:x<0,y=2\pi\mathcal{N}_{k}\}$ if $m_{k}=1$ or $n_{k}=1$. In order to obtain a continuous function we again interpolate between $v_{m_{k+1},n_{k+1}}$ and $v_{m_{k},n_{k}}$. Similarly as before we consider the map $\psi_{k}:(-\infty,0]\to(-\infty,0]$ defined by $v_{m_{k+1},n_{k+1}}(x)=v_{m_{k},n_{k}}(\psi_{k}(x))$. More precisely, denoting by $\phi_{k}$ the function $\phi$ corresponding to $(m_{k},2n_{k})$ and $(m_{k+1},2n_{k+1})$, we have
(3.24) $\begin{split}\psi_{k}(x)=N_{k}\phi_{k}\left(\frac{x}{N_{k+1}}+s_{m_{k+1},n_{k+1}}\right)-N_{k}s_{m_{k},n_{k}}.\end{split}$
Then we define $V:\{z\in\mathbb{C}:\text{Re}\ z\leq 0\}\to\mathbb{C}$ by interpolating between $v_{m_{k+1},n_{k+1}}$ and $v_{m_{k},n_{k}}$ as follows: if $2\pi\mathcal{N}_{k-1}\leq y<2\pi\mathcal{N}_{k}$, say $y=2\pi\mathcal{N}_{k-1}+2\pi N_{k}t$ where $10\leq t<$ 1, then we put
$\begin{split}V(x+iy)=v_{m_{k},n_{k}}((1-t)x+t\psi_{k}(x)+iy)=v_{m_{k},n_{k}}(x+iy+t(\psi_{k}(x)-x)).\end{split}$
This map $V$ is continuous along the horizontal lines $\mathcal{L}^{-}_{k}$ for any $k$ in the left half-plane.
Note that $U$ and $V$ coincide on the imaginary axis. With the definitions of the maps $U$ and $V$, we define the map
(3.25) $\begin{split}G(z)=\left\{\begin{array}[]{ll}U(z),&\text{if}\quad|\arg z|\leq\frac{\pi}{2};\\ V(z),&\text{if}\quad|\arg(-z)|\leq\frac{\pi}{2},\end{array}\right.\end{split}$
which is continuous along the horizontal lines $\mathcal{L}^{+}_{k}$ in the right half-plane and $\mathcal{L}^{-}_{k}$ in the left half-plane for any integer $k$ respectively. In the section below, we will show that the map $G$ defined in (3.25) is indeed quasiregular in the plane and also satisfies the hypothesis of the Teichmüller–Wittich–Belinskii theorem [14, § V.6].
3.3. Estimation of the dilatation
Now we begin to estimate the complex dilatation of the functions $U$, $V$ and $G$ defined in previous section. For a quasiregular map $f$, let
$\begin{split}\mu_{f}(z)=\frac{f_{\bar{z}}(z)}{f_{z}(z)}\ \ \text{and}\ \ K_{f}(z)=\frac{1+|\mu_{f}(z)|}{1-|\mu_{f}(z)|}.\end{split}$
In order to apply the Teichmüller–Wittich–Belinskii theorem, we have to estimate $K_{G}(z)-1$. We note that
$\begin{split}K_{G}(z)-1=\frac{2|\mu_{G}(z)|}{1-|\mu_{G}(z)|}=\frac{2|\mu_{G}(z)|(1+|\mu_{G}(z)|)}{1-|\mu_{G}(z)|^{2}}\leq\frac{4|\mu_{G}(z)|}{1-|\mu_{G}(z)|^{2}}.\end{split}$
By the explicit expressions of $\tilde{g}_{m_{k},n_{k}}$ in (3.20), the arguments for estimating the dilation of $G$ will be similar in either of the three cases (I), (II) and (III). Below we shall first consider the first case (I).
We begin by estimating $K_{U}-1$ in the right half-plane $H^{+}=\{z\in\mathbb{C}:\text{Re}\ z>0\}$. Let $2\pi\mathcal{N}_{k-1}\leq y<2\pi\mathcal{N}_{k}$ so that $10\leq t=(y-2\pi\mathcal{N}_{k-1})/(2\pi N_{k})<$ 1. Then $U(z)=u_{m_{k},n_{k}}(q(z))$ with
(3.26) $\begin{split}q(x+iy)=x+iy+t(\psi_{k}(x)-x)=x+iy+\frac{y-2\pi\mathcal{N}_{k-1}}{2\pi N_{k}}(\psi_{k}(x)-x),\end{split}$
where $\psi_{k}(x)$ takes the form in (3.22). Thus
(3.27) $\begin{split}q_{z}(z)=1+a(z)-ib(z)\ \ \text{and}\ \ q_{\bar{z}}=a(z)+ib(z)\end{split}$
with
(3.28) $\begin{split}a(x+iy)=\frac{t}{2}(\psi_{k}^{\prime}(x)-1)\ \ \text{and}\ \ b(x+iy)=\frac{1}{4\pi N_{k}}(\psi_{k}(x)-x).\end{split}$
Note that if $a(z)<0$, then
$\begin{split}a(z)=\frac{t}{2}(\psi^{\prime}_{k}(x)-1)\geq\frac{1}{2}(\psi^{\prime}_{k}(x)-1)\end{split}$
and thus $01+2a(z)\geq\psi^{\prime}_{k}(z)>$ 0. Then we have $01+2a(z)\geq\min\{1,\psi^{\prime}_{k}(z)\}>$ 0 as well as $01+a(z)>$ 0 in any case.
We deduce from (3.26) that
$\begin{split}|\mu_{U}(z)|^{2}=|\mu_{q}(z)|^{2}=\frac{a(z)^{2}+b(z)^{2}}{(1+a(z))^{2}+b(z)^{2}}\end{split}$
and thus
$\begin{split}K_{U}(z)-1&\leq\frac{4|\mu_{q}(z)|}{1-|\mu_{q}(z)|^{2}}\leq\frac{4\sqrt{(1+a(z))^{2}+b(z)^{2}}\sqrt{a(z)^{2}+b(z)^{2}}}{1+2a(z)}\\ &\leq\frac{4(1+a(z)+|b(z)|)(a(z)+|b(z)|)}{1+2a(z)}\leq\frac{4(1+r(x))r(x)}{\min\{1,\psi_{k}^{\prime}(x)\}},\end{split}$
where
(3.29) $\begin{split}r(x)=|\psi_{k}^{\prime}(x)-1|+|\psi_{k}(x)-x|.\end{split}$
In order to estimate $K_{U}(z)-1$ for $z\in\hat{\mathbf{\Pi}}_{k}^{+}$ defined in (3.21), we recall from Lemma 2.1 that
(3.30) $\begin{split}|\phi_{k}(x)-x|=o(1),\quad x\to\infty.\end{split}$
By the definition of $\psi_{k}(x)$ in (3.22) we have $\psi_{k}(x)-x=l\phi_{k}(x/l+s_{m_{k+1},n_{k+1}})-l(x/l+s_{m_{k+1},n_{k+1}})+l(s_{m_{k+1},n_{k+1}}-s_{m_{k},n_{k}})$ and, together with the estimate in (3.30), thus
$\begin{split}|\psi_{k}(x)-x|=O(1),\quad x\to\infty.\end{split}$
Further, by Lemma 2.1 we also have
$\begin{split}|\psi^{\prime}_{k}(x)-1|=o(1),\quad x\to\infty.\end{split}$
The last two equations imply that the functions $r(x)$ defined by (3.29) satisfies $r(x)=O(1)$ for $z\in\hat{\mathbf{\Pi}}_{k}^{+}$. Moreover, we note that the affine mapping $\chi_{m_{k},n_{k}}:H^{+}\to H^{+}$ defined in (3.18) satisfies $\chi_{m_{k},n_{k}}(z)=x/l+iy/N_{k}$ and so $|\mu_{\chi_{m_{k},n_{k}}}|=(N_{k}-l)/(N_{k}+l)<1$. Thus $U$ is quasiregular in the right half-plane $H^{+}$ and
$\begin{split}K_{U}(z)-1=O(1),\quad x\to\infty\end{split}$
for $z\in\hat{\mathbf{\Pi}}_{k}^{+}$. We denote
[…]
and
[…]
Let $\varepsilon>0$ be small. Define $\gamma$ as follows: let $\gamma=\varepsilon$ when $\lambda_{2}=0$ and $\gamma=\lambda_{2}$ when $\lambda_{2}>0$. Then, together with the estimates in (3.9) and (3.10), we have
$\begin{split}\int_{S_{1,k_{i}}}\frac{dxdy}{x^{2}+y^{2}}\leq\int_{S_{1,k_{i}}}\frac{dxdy}{x^{2}+4\pi^{2}(k_{i}-1)^{2}}\leq 2\pi N_{k_{i}}\int_{0}^{\infty}\frac{dx}{x^{2}+4\pi^{2}\mathcal{N}_{k_{i}-1}^{2}}=\frac{\pi N_{k_{i}}}{2\mathcal{N}_{k_{i}-1}}.\end{split}$
By the choice of the two sequence of $(m_{k})$ and $(n_{k})$ we easily see that
(3.31) $\begin{split}\mathcal{N}_{k}\sim lk\quad\text{as}\quad k\to\infty.\end{split}$
This yields
$\begin{split}\int_{S_{1,k_{i}}}\frac{dxdy}{x^{2}+y^{2}}\leq\frac{\pi}{c(2\pi i)^{1/\gamma}}\end{split}$
for $i\geq 1$ and some positive constant $c$. Thus
[…]
for all $i\geq 1$ and some constant $c_{1}$. Similarly, we also have
[…]
for all $i\geq 1$ and some constant $c_{2}$. Moreover, together with the estimates in (3.11) and (3.12), we can also deduce that
[…]
for all $j\geq 1$ and some constant $c_{3}$, and that
[…]
for all $j\geq 1$ and some constant $c_{4}$. Note that $U$ is meromorphic in $H^{+}$ apart from the half-strips $S_{1,k_{i}-1}$, $S_{1,k_{i}}$, $S_{1,k_{j}-1}$ and $S_{1,k_{j}}$. Then the last four inequalities (3.32)-(3.35) imply that
[…]
The estimate of $K_{V}(z)-1$ in the left half-plane $H^{-}=\{z\in\mathbb{C}:\text{Re}\ z<0\}$ is similar. Now we have $V(z)=V_{m_{k},n_{k}}(q(z))$ where $q(z)$ has the same form as in (3.26). Moreover, (3.27) and (3.28) both hold for the same $a(x+iy)$ and $b(x+iy)$, but $\psi_{k}(x)$ taking the form in (3.24). Instead of (3.28) we obtain
$\begin{split}K_{V}(z)-1&\leq\frac{4(1+r(x))r(x)}{\min\{1,\psi_{k}^{\prime}(x)\}}\end{split}$
with
(3.37) $\begin{split}r(x)=\left|\psi^{\prime}_{k}(x)-1\right|+\frac{1}{N_{k}}\left|\psi_{k}(x)-x\right|.\end{split}$
Now, by the definition of $\psi_{k}(x)$ in (3.24), we have
$\begin{split}&\frac{1}{N_{k}}\left|\psi_{k}(x)-x\right|=\left|\phi_{k}\left(\frac{x}{N_{k+1}}+s_{m_{k+1},n_{k+1}}\right)-\frac{x}{N_{k}}-s_{m_{k},n_{k}}\right|\\ =&\ \left|\phi_{k}\left(\frac{x}{N_{k+1}}+s_{m_{k+1},n_{k+1}}\right)-\frac{N_{k+1}}{N_{k}}\left(\frac{x}{N_{k+1}}+s_{m_{k+1},n_{k+1}}\right)+\frac{N_{k+1}}{N_{k}}s_{m_{k+1},n_{k+1}}-s_{m_{k},n_{k}}\right|\end{split}$
and
$\begin{split}\left|\psi^{\prime}_{k}(x)-1\right|=\frac{N_{k}}{N_{k+1}}\left|\phi^{\prime}_{k}\left(\frac{x}{N_{k+1}}+s_{m_{k+1},n_{k+1}}\right)-\frac{N_{k+1}}{N_{k}}\right|.\end{split}$
We thus have by Lemma 2.1 that
$\begin{split}&\frac{1}{N_{k}}\left|\psi_{k}(x)-x\right|=O(1),\quad x\to-\infty\end{split}$
and also that
$\begin{split}&\left|\psi^{\prime}_{k}(x)-1\right|=o(1),\quad x\to-\infty.\end{split}$
The last two inequalities imply that the function $r(x)$ defined in (3.37) satisfies $r(x)=O(1)$ for $z\in\hat{\mathbf{\Pi}}_{k}^{-}$ defined in (3.23).
[…]
and
[…]
Then by the same arguments as the estimate for $K_{U}-1$ in the right-half plane, we also conclude that $V$ is quasiregular in the left half-plane and that
$\begin{split}K_{V}(z)-1=O(1),\quad x\to-\infty\end{split}$
for $z\in T_{1,k_{i}}$, $z\in T_{1,k_{i}+1}$, $z\in T_{1,k_{j}}$ and $z\in T_{1,k_{j}+1}$. This finally implies that
[…]
The two inequalities in (3.36) and (3.38) shows that, if $r>0$, then
[…]
For the case (II), since $l=3$ and since the sequence $(n_{k})$ are chosen in the way that $n_{k}=1$ for all $k\geq 4$, we also have we have $N_{k}\leq 4$ for all $k\geq 4$ and we only need to estimate the complex dilatation of $U$ in the stripes $S_{1,k_{i}-1}$ and $S_{1,k_{i}}$ of $\hat{\mathbf{\Pi}}_{k}^{+}$ and the dilation of $V$ in the stirps $T_{1,k_{i}-1}$ and $T_{1,k_{i}}$ of $\hat{\mathbf{\Pi}}_{k}^{-}$ as well as the complex dilatation of $U$ in the stripes $\hat{\mathbf{\Pi}}_{k}^{+}$ for the case $k=1,2,3$ to conclude that the inequality in (3.39) also holds.
For the case (III), since $l=4$ and since the two sequences $(m_{k})$ and $(n_{k})$ are chosen in the way that $m_{k}=n_{k}=1$ for all $k\geq 4$, we have $N_{k}=4$ for all $k\geq 4$. Thus we only need to estimate the complex dilatation of $U$ in the stripes $\hat{\mathbf{\Pi}}_{k}^{+}$ for the case $k=1,2,3$ to conclude that the inequality in (3.39) also holds.
Therefore, in either of the three cases (I), (II) and (III), the function $G$ defined in (3.25) satisfies the hypothesis of the Teichmüller–Wittich–Belinskii theorem. This theorem, together with the existence theorem for quasiconformal mappings, yields that there exists a quasiconformal homeomorphism $\tau:\mathbb{C}\to\mathbb{C}$ and a meromorphic function $F$ such that
(3.40) $\begin{split}G(z)=F(\tau(z))\quad\text{and}\quad\tau(z)\sim z\quad\text{as}\quad z\to\infty.\end{split}$
Then $F=G\circ\tau^{-1}$ is a locally univalent meromorphic function. Define $E=F/F^{\prime}$. Then $E$ is a Bank–Laine function. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A$.
3.4. Proof for the case $n=1$
We first prove Theorem 1.4 for the case $n=1$. We shall consider the three cases (I), (II) and (III) together. We will use similar arguments as in the proof of Theorem 1.3 to estimate the numbers of poles and zeros of $F$ respectively and also to determine the orders of $E$ and $A$. Also, for two positive quantities $S_{1}(r)$ and $S_{2}(r)$, we use the notation $S_{1}(r)\asymp S_{2}(r)$, provided that there are two positive constants $c_{1}$ and $c_{2}$ such that $c_{2}S_{2}(r)\leq S_{1}(r)\leq c_{1}S_{2}(r)$ for all large $r$.
We first estimate the number of poles of $F$. Let $r>0$ and choose $k\in\mathbb{N}$ such that $2\pi\mathcal{N}_{k-1}<r\leq 2\pi\mathcal{N}_{k}$. Note that $\mathcal{N}_{k}\sim lk$ as $k\to\infty$ by (3.31). By the construction of $(m_{k})$ together with the definition of $\chi_{m,n}$ in (3.18), we have
(3.41) $\left\{\begin{array}[]{ll}\sum_{i=1}^{k}m_{i}\leq n(r,\infty,U)\leq 2\sum_{i=1}^{k}m_{i},&\\ \sum_{i=1}^{k}m_{i}\leq n(r,\infty,V)\leq 2\sum_{i=1}^{k}m_{i}.&\end{array}\right.$
By the choice of $k_{i}$ in (3.6) or (3.7) in the proof of Lemma 3.1, we easily see that $\sum_{i=1}^{k}m_{i}\sim(2\pi)^{-1}(\log(2\pi k))^{2}$ as $k\to\infty$ when $\lambda_{1}=0$ and that $\sum_{i=1}^{k}m_{i}\sim(2\pi)^{\lambda_{1}-1}k^{\lambda_{1}}$ as $k\to\infty$ when $0<\lambda_{1}\leq 1$. Together with these two relations, the two double inequalities in (3.41) imply that $n(r,\infty,G)\asymp(\log r)^{2}$ when $\lambda_{1}=0$ and that $n(r,\infty,G)\asymp r^{\lambda_{1}}$ when $0<\lambda_{1}\leq 1$. Then the relations in (3.40) imply that $n(r,\infty,F)\asymp(\log r)^{2}$ when $\lambda_{1}=0$ and that $n(r,\infty,F)\asymp r^{\lambda_{1}}$ when $0<\lambda_{1}\leq 1$. Thus $\lambda(f_{1})=\lambda_{1}$ in all three cases.
The estimate for the number of zeros of $F$ is similar. Again, let $r>0$ and choose $k\in\mathbb{N}$ such that $2\pi\mathcal{N}_{k-1}<r\leq 2\pi\mathcal{N}_{k}$. By the construction of the sequence $(n_{k})$ together with the definition of $\chi_{m,n}$ in (3.18), we have
(3.42) $\left\{\begin{array}[]{ll}2\sum_{j=1}^{k}n_{j}\leq n(r,0,U)\leq 4\sum_{i=j}^{k}n_{j},&\\ 2\sum_{j=1}^{k}n_{j}\leq n(r,0,V)\leq 4\sum_{j=1}^{k}n_{j}.&\end{array}\right.$
By the same arguments as before, we have $2\sum_{j=1}^{k}n_{j}\sim(2\pi)^{-1}(\log(2\pi k))^{2}$ as $k\to\infty$ when $\lambda_{2}=0$ and that $2\sum_{j=1}^{k}n_{j}\sim(2\pi)^{\lambda_{2}-1}k^{\lambda_{2}}$ as $k\to\infty$ when $0<\lambda_{2}\leq 1$. Together with these two relations, the two double inequalities in (3.42) imply that $n(r,0,G)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and that $n(r,0,G)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Now the relations in (3.40) imply $n(r,0,F)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and that $n(r,0,F)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Thus $\lambda(f_{2})=\lambda_{2}$ in all three cases.
Next, let $\varepsilon>0$ be a small constant. Since $m,n\leq 1$ for all $k$, there is a set $\Omega_{1}$ of finite linear measure such that the functions $P_{n}(e^{z})$ and $Q_{m}(e^{z})$ satisfy
(3.43) $\left\{\begin{array}[]{ll}\exp(-(2n+\varepsilon)r)\leq\left|P_{n}(e^{z})\right|\leq\exp((2n+\varepsilon)r),&\\ \exp(-(m+\varepsilon)r)\leq\left|Q_{m}(e^{z})\right|\leq\exp((m+\varepsilon)r)&\end{array}\right.$
both hold for all $z\in\mathbb{C}$ such that $|z|\not\in\Omega_{1}$. By the choice of the two sequences $(m_{k})$ and $(n_{k})$ we may suppose that the two double inequalities in (3.43) hold for all pairs $(m_{k},2n_{k})$. This implies that
(3.44) $\begin{split}|g_{m,n}(z)|\leq\exp((m+\varepsilon)r+(2n+\varepsilon)r)\left|\exp(e^{z})\right|\end{split}$
for all $z\in\mathbb{C}$ such that $|z|\not\in\Omega_{1}$ and all pairs $(m,2n)$. For simplicity, we denote $L=m+2n$ and $L_{k}=m_{k}+2n_{k}$. Clearly this implies that there is a set $\Omega_{2}$ of finite linear measure such that
(3.45) $\left\{\begin{array}[]{ll}|v_{m,n}(z)|\leq\exp((L+2\varepsilon)r/N)\left|\exp(e^{z/N})\right|&\text{for}\quad z\in H^{-}\quad\text{and}\quad|z|\not\in\Omega_{2},\\ |u_{m,n}(z)|\leq\exp((L+2\varepsilon)r/l)\left|\exp(e^{z/l})\right|&\text{for}\quad z\in H^{+}\quad\text{and}\quad|z|\not\in\Omega_{2}\end{array}\right.$
for all pairs $(m,2n)$.
Let $z=x+iy\in H^{-}$ with $\text{Im}\ z\geq 0$ and $k\in\mathbb{N}$ with $2\pi\mathcal{N}_{k-1}\leq\text{Im}\ z<2\pi\mathcal{N}_{k}$. With $t=(y-2\pi\mathcal{N}_{k-1})/{2\pi N_{k}}$ we have $10\leq t<$ 1 and
(3.46) $\begin{split}|V(x+iy)|=|v_{m_{k},n_{k}}((1-t)x+t\psi_{k}(x)+iy)|\leq\exp((L_{k}+3\varepsilon)r/N_{k}),\end{split}$
where $\psi_{k}(x)$ is defined in (3.24). Thus we have
(3.47) $\begin{split}|V(z)|\leq\exp((L+3\varepsilon)r)\end{split}$
for $z\in H^{-}$ and $|z|\not\in\Omega_{2}$. For $z\in H^{+}$ and $|z|\not\in\Omega_{2}$, by the inequality in (3.44) we have
$\begin{split}|g_{m,n}(z)|\leq\exp((L+2\varepsilon)r/l)\exp\left(e^{(\text{Re}\ z)/l}\right).\end{split}$
Again, assuming that $\text{Im}\ z\geq 0$, we choose $k\in\mathbb{N}$ such that $2\pi\mathcal{N}_{k-1}\leq\text{Im}\ z<2\pi\mathcal{N}_{k}$. Then, with $U(z)=u_{m_{k},n_{k}}(q(z))=g_{m_{k},n_{k}}(q(z)+s_{m_{k},n_{k}})$ where $q(z)$ is defined by (3.26), we have
(3.48) $\begin{split}|U(z)|\leq\exp((L_{k}+2\varepsilon)(\text{Re}(q(z))/l+s_{m_{k},n_{k}})\exp\left(e^{\text{Re}(q(z))/l+s_{m_{k},n_{k}}}\right).\end{split}$
Since $m_{k}$ and $n_{k}$ are either 00 or $1$ and $s_{m_{k},n_{k}}=O(1)$, we deduce from (3.26) and (3.48) that
$\begin{split}\log|U(z)|&\leq(L_{k}+2\varepsilon)\text{Re}(q(z))/l+s_{m_{k},n_{k}}+\exp(\text{Re}(q(z))/l+s_{m_{k},n_{k}})\\ &\leq\exp((1+o(1))|z|/l)+O(1)=\exp((1+o(1))|z|/l),\quad z\to\infty\quad\text{in}\quad H^{+}.\end{split}$
Together with (3.47) we conclude that
$\begin{split}\log|G(z)|\leq\exp((1+o(1))|z|/l)\end{split}$
as $|z|\to\infty$ outside a set $\Omega_{3}$ of finite linear measure. Hence (3.40) yields that
$\begin{split}\log|F(z)|\leq\exp((1+o(1))|z|/l)\end{split}$
as $|z|\to\infty$ outside a set $\Omega_{4}$ of finite linear measure. The lemma on the logarithmic derivative now implies that $E=F/F^{\prime}$ satisfies
(3.49) $\begin{split}m\left(r,\frac{1}{E}\right)=O(r)\end{split}$
as $|z|\to\infty$ outside a set of finite linear measure. Moreover, since all zeros of $E$ are simple and since $\lambda_{1}\leq\lambda_{2}$, the previous estimates on the poles and zeros of $F$ imply that, when $\lambda_{2}=0$,
(3.50) $\begin{split}N(r,0,E)\leq c_{1}(\log r)^{3}\end{split}$
for some positive constant $c_{1}$ and all large $r$ or, when $0<\lambda_{2}\leq 1$,
(3.51) $\begin{split}c_{3}r^{\lambda_{2}}\leq N(r,0,E)\leq c_{2}r^{\lambda_{2}}\end{split}$
for two positive constants $c_{2},c_{3}$ and all large $r$. By the estimate in (3.49) together with the estimates in (3.50) or (3.51), the first main theorem implies that $T(r,E)=T(r,1/E)+O(1)=O(r)$. Thus $\lambda(E)\leq\rho(E)\leq 1$. In particular, if $\lambda_{2}=1$, then the estimate in (3.51) shows that $\lambda(E)=\rho(E)=1$. If $\lambda_{2}<1$, then the classical Hadamard factorization theorem together with the lemma on the logarithmic derivative easily yields that $T(r,E^{\prime}/E)=o(1)T(r,E)$ and $T(r,E^{\prime\prime}/E)=o(1)T(r,E)$ where $r\to\infty$. Then the relation in (1.8) implies that $T(r,A)=T(r,1/E)+o(1)T(r,E)=T(r,E)+o(1)T(r,E)$. Since an independent estimate below which only uses the fact that $E$ is of finite order yields $\rho(A)=1$, we also have $\rho(E)=1$.
Now, since $E$ is of finite order, it follows by the relation (1.8) together with the lemma on the logarithmic derivative that
$\begin{split}m(r,A)=m\left(r,\frac{1}{E}\right)+O(\log r)=O(r).\end{split}$
Thus the first main theorem implies that $\rho(A)\leq 1$. On the other hand, we note that $g_{1,1}$ is analytic in $\hat{\mathbf{\Pi}}_{1}^{+}$ as well as in $\check{\mathbf{\Pi}}_{-1}^{+}$. We see from the definition of $\tilde{g}_{m,n}$ in (3.18) that $\chi_{1,1}(z)=x/l+iy$ in the right half-plane and commutes with the complex conjugation. Together with the function $\tau$ in (3.40), we have $|(\chi_{1,1}\circ\tau^{-1})(z)|\leq 2|z|/l$ for $z$ such that $|z|$ is large. Denote $\tilde{\chi}_{1,1}=\chi_{1,1}+s_{1,1}$. Then by the same arguments as in the proof of Theorem 1.3 we may choose a curve in $(\tau\circ\tilde{\chi}_{1,1}^{-1})(\hat{\mathbf{\Pi}}_{1}^{+}\cup\check{\mathbf{\Pi}}_{-1}^{+})$, say $\mathcal{L}^{+}$, such that $(\tilde{\chi}_{1,1}\circ\tau^{-1})(\mathcal{L}^{+})$ is the positive real axis and also that $\tau\circ\tilde{\chi}_{1,1}^{-1}$ satisfies $|(\tau\circ\tilde{\chi}_{1,1}^{-1})^{\prime}(z)|\leq l|2z|$ for all $z\in(\tilde{\chi}_{1,1}\circ\tau^{-1})(\mathcal{L}^{+})$ such that $|z|$ is large. Then, by the relation in (1.8) as well as the definitions of $u_{m,n}$ in Subsection 3.2 and the relation $F=G\circ\tau^{-1}$, we easily see that $\log|A|\geq c_{4}|z|$ for some positive constant $c_{4}$ and all $z\in\mathcal{L}^{+}$ such that $|z|$ is large, implying that $\rho(A)\geq 1$. Hence $\rho(A)=1$. This proves Theorem 1.4 for the case $n=1$.
3.5. Proof for the case $n\geq 2$
In this subsection, we show how to extend the results in the case $n=1$ to the case $n\geq 2$ for any positive integer $n$ with the same process as in [5, Section 5]. For two numbers $\tilde{\lambda}_{1},\tilde{\lambda}_{2}\in[0,n]$ such that $\Lambda_{1}\leq\Lambda_{2}$, we let $\lambda_{1}=\Lambda_{1}/n$ and $\lambda_{2}=\Lambda_{2}/n$. Then, we choose two sequences $(m_{k})$ and $(n_{k})$ associated with $\lambda_{1}$ and $\lambda_{2}$ in the same way as in Subsection 3.2. Note that $m_{k}=n_{k}=0$, $k=1,2,3$, for the two sequences $(m_{k})$ and $(n_{k})$ in any case.
Let $\chi_{1,1}:\mathbb{C}\to\mathbb{C}$ be such that $\chi_{1,1}(z)=x/l+iy$ when $x\geq 0$ and $\chi_{1,1}(z)=x+iy$ when $x\leq 0$. Here $l=1$ when $\lambda_{1}\leq\lambda_{2}<1$ and $l=3$ when $\lambda_{1}<\lambda_{2}=1$ and $l=4$ when $\lambda_{1}=\lambda_{2}=1$. Then we choose the numbers $\rho_{0}$ and $\sigma_{0}$ in [5, Section 5] to be $\rho_{0}=\sigma_{0}=1$ and the quasiconformal mapping $Q$ there to be $Q(z)=\chi_{1,1}(z)$.
We define a similar function $G_{0}$ as in previous section. Denote by $J^{+}$ and $J^{-}$ the preimages of the half-strips $\{z\in\mathbb{C}:\text{Re}\ z\geq 0,0\leq\text{Im}\ z\leq 2\pi\}$ and $\{z\in\mathbb{C}:\text{Re}\ z\leq 0,0\leq\text{Im}\ z\leq 2\pi\}$ under $z\mapsto\chi_{1,1}(z)$ respectively. With $s_{0}=\log\log 2$, we then have
(3.52) $G_{0}(z)=\left\{\begin{array}[]{ll}\exp\exp(\chi_{1,1}(z)+s_{0}),&\text{if}\quad z\in J^{+},\\ \exp\exp(z+s_{0}),&\text{if}\quad z\in J^{-}.\end{array}\right.$
We note that $J:=J^{+}\cup J^{-}$ is bounded by the real axis and a line in the upper half-plane which, as the real axis, is mapped to $(1,\infty)$ by $G_{0}$. In addition to $G_{0}$, we will also consider a modification $G_{1}$ of $G_{0}$ defined as follows. Let $U$ and $V$ be as in Subsection 3.2. For $\text{Re}\ z\geq 0$ we define
$U_{1}(z)=\left\{\begin{array}[]{lll}U(z-\pi i),&\text{if}\quad\text{Im}\ z\geq\pi,\\ \exp(-\exp(z+s_{0}),&\text{if}\quad-\pi<\text{Im}\ z<\pi,\\ U(z+\pi i),&\text{if}\quad\text{Im}\ z\leq-\pi\end{array}\right.$
and for $\text{Re}\ z<0$ we define
$V_{1}(z)=\left\{\begin{array}[]{lll}V(z-\pi i),&\text{if}\quad\text{Im}\ z\geq\pi,\\ \exp(-\exp(z+s_{0}),&\text{if}\quad-\pi<\text{Im}\ z<\pi,\\ V(z+\pi i),&\text{if}\quad\text{Im}\ z\leq-\pi.\end{array}\right.$
The map $G_{1}$ is then obtained by gluing $U_{1}$ and $V_{1}$ in the same way in which $U$ and $V$ were glued to obtain $G_{0}$. Thus we obtain
(3.53) $G_{1}(z)=\left\{\begin{array}[]{ll}\exp(-\exp(\chi_{1,1}(z)+s_{0})),&\text{if}\quad z\in J^{+},\\ \exp(-\exp(z+s_{0})),&\text{if}\quad z\in J^{-}.\end{array}\right.$
It follows that $G_{1}(z)=1/G_{0}(z)$ for $z\in J$.
For $j=1,\cdots,2n$ we put
$\begin{split}\Sigma_{j}=\left\{z\in\mathbb{C}:(j-1)\frac{\pi}{n}<\arg z<j\frac{\pi}{n}\right\}.\end{split}$
Here we have used the notation $n$ instead of the $N$ in [4, Section 5]). Let $H=\{z\in\mathbb{C}:\text{Im}\ z>0\}$ be the upper half-plane and $L=H/J$. Denote by $\bar{L}$ the reflection of $L$ on the real axis; that is, $\bar{L}=\{z\in\mathbb{C}:\bar{z}\in L\}$. Denote by $D_{j}$ the preimage of $L$ or $\bar{L}$ under the map $z\mapsto z^{n}$ in $\Sigma_{j}$ for $j=1,\cdots,2n$. We will define the quasiregular map $G:\mathbb{C}\to\mathbb{C}$ which satisfies $G(z)=G_{0}(z^{n})$ for $z\in D_{j}$ if $12\leq j\leq 2N-$ 1 and $G(z)=G_{1}(z^{n})$ for $z\in D_{1}$ and $z\in D_{2n}$. In the remaining part of the plane we will define $G$ by a suitable interpolation which will require only (3.52) and (3.53). In order to do this, we will first define $G$ in a neighborhood of $\infty$. Thus, for a suitable $r_{0}>0$, we have to define $G$ in certain neighborhoods of the rays $\{z\in\mathbb{C}:\arg z=j\pi/n,|z|>r_{0}\}$. If $j\not=1$ and $j\not=2n-1$, then $G$ is defined by the same expression in the domains $D_{j}$ and $D_{j+1}$ adjacent to the ray. (Here we have put $D_{2n+1}=D_{1}$. In similar expressions the index $j$ will also be taken modulo $2n$.) For $j=1$ and $j=2n-1$ the interpolation argument is more elaborate. Finally, the definition of $G$ can be extended to the remains.
By following similar process as in [4, Section 5]) we may finally obtain a locally univalent meromorphic function $F=G\circ\tau^{-1}$ for a quasiconformal mapping $\tau$ in the plane satisfying $\tau(z)\sim z$ as $z\to\infty$. We shall not repeat the details. Then $E=F/F^{\prime}$ is a Bank–Laine function and $E=f_{1}f_{2}$ for two normalized solutions of the second order linear differential equation (1.1) with an entire function $A$.
As in the case $n=1$, we may estimate the numbers of poles or zeros of $G_{0}(z^{n})$ in the sectors $D_{2},D_{3},\cdots,D_{2n-1}$ and that of $G_{1}(z^{n})$ in the sectors $D_{1},D_{2n}$, respectively. Since $G$ has no poles or zeros in the region $\mathbb{C}-\sum_{j=1}^{2n}D_{j}$, this finally yields that $n(r,\infty,F)\asymp(\log r)^{2}$ when $\Lambda_{1}=0$ and that $n(r,\infty,F)\asymp r^{\Lambda_{1}}$ when $0<\Lambda_{1}\leq n$ and also that $n(r,0,F)\asymp(\log r)^{2}$ when $\Lambda_{2}=0$ and that $n(r,0,F)\asymp r^{\Lambda_{2}}$ when $0<\Lambda_{2}\leq n$. Moreover, by the construction of $G$, we may estimate $\log|G|$ from above in a similar way as in the case $n=1$ to obtain that $\log|F|\leq\exp((1+o(1))|z|^{n}/l)$ as $z\to\infty$, which implies that $\rho(E)\leq n$ and also $\rho(A)\leq n$. On the other hand, by the definition of $G$ in the remaining parts $\mathbb{C}-\sum_{j=1}^{2n}D_{j}$, we may estimate $|A|$ from below along certain curve in similar way as in the case $n=1$. For example, letting $\eta:\mathbb{C}\to\mathbb{C}$ be the map such that $\eta(z)=z^{n}$, we may choose a curve in the closure of $(\tau\circ\eta^{-1}\circ\tilde{\chi}_{1,1}^{-1})(\eta^{-1}{(\Sigma_{2})}\cup\eta^{-1}(\Sigma_{3}))$, say $Y_{2}^{\prime}$, so that $(\tilde{\chi}_{1,1}\circ\eta\circ\tau^{-1})(Y_{2}^{\prime})$ is the positive real axis. For a point $z$ in the real axis, we may use the Cauchy formula as in the proof of Theorem 1.3 to show that $|(\tau\circ\eta^{-1}\circ\tilde{\chi}_{1,1}^{-1})^{\prime}(z)|\leq l|2z|^{1/n}$ when $|z|$ is large, which then yields that $\rho(A)\geq n$. Finally we conclude that $\rho(E)=\rho(A)=n$. We shall not repeat the details here. This completes the proof of Theorem 1.4.
4. Proof of Theorem 1.6
4.1. Preliminary lemmas
We need to prepare some preliminary lemmas which are slight modifications of those in [19, Section 2]. We shall first provide the corrected versions of two lemmas in [19, Lemma 2.1 and Lemma 2.2]. Since the construction of the Bank–Laines $E$ in the case $\lambda(E)=1$ has been included in Theorem 1.4, here we only consider the case where $\lambda(E)>1$.
Lemma 4.1.
Let $\gamma\in(1,\infty)$ and $\delta\in[0,1]$ be two numbers. Then there exists a nonnegative sequence $(m_{k})$ of integers, with $m_{k}=0$ for $k=1,2,3$, such that the function $\mathfrak{h}:[0,\infty)\to[0,\infty)$ which satisfies $\mathfrak{h}(0)=0$ and which is linear on each interval $[2\pi(k-1),2\pi k]$ and has slopes $m_{k}$ satisfies either
  • (1)
    when $0\leq\delta\gamma\leq 1$, there is a subsequence $(m_{k_{i}})$ such that $m_{k_{i}}=1$ for $k\geq k_{1}$ and $m_{k}=0$ for other $k$ such that
    $\begin{split}\mathfrak{h}(x)=x^{\delta\gamma}+O(1),\quad x\to\infty\end{split}$
    and, in particular, $m_{k}=0$ for all $k\geq k_{1}$ when $\delta\gamma=0$ and $m_{k}=1$ for all $k\geq k_{1}$ when $\delta\gamma=1$ and $k_{i}\geq ci^{1/\delta\gamma}$ for a fixed positive constant $c$ when $10<\delta\gamma<$ 1; or
  • (2)
    when $\delta\gamma>1$, $(m_{k})$ is increasing such that $m_{k+1}-m_{k}$ is an odd integer when $m_{k+1}>m_{k}$ such that
    $\begin{split}\mathfrak{h}(x)=(\log(x+2\pi))^{-2}x^{\delta\gamma}+O\left((\log(x+2\pi))^{-1}x^{\delta\gamma-2}\right)+O(1),\quad x\to\infty\end{split}$
    and
    $\begin{split}m_{k}=\delta\gamma(\log(k+2\pi))^{-2}(2\pi k)^{\delta\gamma-1}+O\left((\log(k+2\pi))^{-2}k^{\delta\gamma-2}\right)+O(1),\quad k\to\infty.\end{split}$
Lemma 4.2.
Let $\gamma\in(1,\infty)$ be the constant in Lemma 4.1 and $(m_{k})$ be the sequence there. Then there exists a sequence $(n_{k})$ of nonnegative integers, with $n_{k}=0$ for $k=1,2,3$, such that the function $h:[0,\infty)\to[0,\infty)$ which satisfies $h(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ and has slopes $N_{k}=m_{k}+2n_{k}+1$ there satisfies
$\begin{split}h(x)=x^{\gamma}+O(x^{\gamma-2})+O(1),\quad x\to\infty\end{split}$
and
(4.1) $\begin{split}N_{k}=m_{k}+2n_{k}+1=\gamma(2\pi k)^{\gamma-1}+O\left(k^{\gamma-2}\right)+O(1),\quad k\to\infty.\end{split}$
Moreover, the function $g(x)$ defined on $[0,\infty)$ by $h(g(x))=x^{\gamma}$ satisfies
(4.2) $\begin{split}g(x)=x+O(x^{-1})+O(x^{1-\gamma}),\quad x\to\infty\end{split}$
and
(4.3) $\begin{split}g^{\prime}(x)=1+O(x^{-1})+O(x^{1-\gamma}),\quad x\to\infty,\end{split}$
where $g^{\prime}$ denotes either the left or right derivative of $g$.
For two numbers $\gamma\in(1,\infty)$ and $\delta\in[0,1]$, we choose $(m_{k})$ to be the sequence constructed in Lemma 4.1. Note that $(m_{k})$ can be modified to have the properties in Lemma 3.1 in the case $\delta=0$. Then we may follow the proof in [19, Lemma 2.1] to find a suitable sequence $(n_{k})$ such that the sequence $(N_{k})$ where $N_{k}=m_{k}+2n_{k}+1$ has the desired properties in Lemma 4.2 for the functions $h$ and $g$ there. Moreover, the assumptions that $m_{k}=0$ for $k=1,2,3$ in Lemma 4.1 and that $n_{k}=0$ for $k=1,2,3$ in Lemma 4.2 can be satisfied by slightly modifying the original proofs in [19].
In the original version of Lemma 4.2 (i.e.,[19, Lemma 2.1]), we have chosen the slopes of $h$ in the interval $[2\pi(k-1),2\pi k]$ to be $12n_{k}+$ 1. However, this then yields the function $g$ defined by $h(g(x))+\mathfrak{h}(g(x))=x^{\gamma}$ there deviates from the positive real axis too far away in the sense that $g(x)-x$ are unbounded as $x\to\infty$. For such a function $g$, the corresponding function $Q$ in the definition of $G$ in [19, Subsection 2.2] may not be quasiconformal in the plane. Now, by Lemma 4.2, we have the relations in (4.2) and (4.3) such that $g(x)-x$, as well as $g^{\prime}(x)-1$, is bounded as $x\to\infty$. With the corrected versions of Lemma 4.1 and Lemma 4.2, the proof for Theorem 1.2 is complete.
Below we begin to prove Theorem 1.6 for the case $\rho\in(1/2,1)$. Let $\gamma\in(1,\infty)$ and $\delta\in[0,1]$. For a pair $(m,2n)$ of two nonnegative integers, we shall first modify $g_{m,n}$ in (1.5) slightly to define
(4.4) $\begin{split}\hat{g}_{m,n}(z)=\frac{g_{m,n}(z)+1}{2}.\end{split}$
Actually we shall restrict the functions in (4.4) to the region $\{z=x+iy,0\leq y\leq 4\pi\}$ of the upper half-plane and to the rest of the plane we shall restrict the functions
(4.5) $\begin{split}\check{g}_{m,n}(z)=g_{m,n}(z).\end{split}$
Thus we may write
(4.6) $\begin{split}\bar{g}_{m,n}(z)=(\bar{\chi}_{m,n}\circ g_{m,n})(z)\end{split}$
for an affine mapping $\bar{\chi}:\mathbb{C}\to\mathbb{C}$ such that either $\bar{\chi}_{m,n}(z)=(z+1)/2$ or $\bar{\chi}_{m,n}(z)=z$. We choose the sequence $(m_{k})$ and $(n_{k})$ constructed in Lemma 4.1 and Lemma 4.2 respectively. The form of $\bar{\chi}$ defined in (4.6) will depend on the pair $(m_{k},2n_{k})$ of two nonnegative integers $m_{k}$ and $n_{k}$. The assumptions that $m_{k}=n_{k}=0$ for $k=1,2,3$ make it convenient to extend the construction of the entire function $A$ in the case $\rho(A)=\rho\in(1/2,1)$ to the case $\rho(A)=\rho\in(n/2,n)$ for any positive integer $n$ as in [5, Section 5].
Let $m,n,\mathfrak{m},\mathfrak{n}\in\mathbb{N}$ and put $N=m+2n+1$ and $M=\mathfrak{m}+2\mathfrak{n}+1$. By the definition of $\bar{\chi}_{m,n}$ in (4.6), for each pair $(m,2n)$ of two nonnegative integers $m$ and $n$, we shall consider the function $\bar{\phi}:\mathbb{R}\to\mathbb{R}$ defined by
(4.7) $\begin{split}(\bar{\chi}_{\mathfrak{m},\mathfrak{n}}\circ g_{\mathfrak{m},\mathfrak{n}})(x)=\bar{g}_{\mathfrak{m},\mathfrak{n}}(x)=(\bar{g}_{m,n}\circ\bar{\phi})(x)=(\bar{\chi}_{m,n}\circ g_{m,n}\circ\bar{\phi})(x).\end{split}$
We will consider the functions $\bar{\phi}$ for the case that $m=m_{k}$, $n=n_{k}$, $\mathfrak{m}=m_{k+1}$ and $\mathfrak{n}=n_{k+1}$, that is, $M=N_{k+1}$ and $N=N_{k}$. Thus we will consider the asymptotic behaviors of $\bar{\phi}_{k}$ as $k\to\infty$, but in order to simplify the formulas in most situations below we suppress the dependence of $\bar{\phi}$ from $m,n$ and $\mathfrak{m},\mathfrak{n}$ from the notation.
We first fix two integers $M$ and $N$ for simplicity and suppose that $\bar{\chi}_{\mathfrak{m},\mathfrak{n}}(z)=(z+1)/2$ and $\bar{\chi}_{m,n}(z)=z$. Let $K=M/N$. By following exactly the same arguments as in the proof of Lemma 2.1, it is easy to check that
(4.8) $\begin{split}\bar{\phi}(x)&=x+O(e^{-x/2}),\ \ \ \ \ \ \ \ \ \ \ \ \ \ \bar{\phi}^{\prime}(x)=1+O(e^{-x/2}),\quad x\to\infty,\\ \bar{\phi}(x)&=Kx+c+O(e^{-\delta|x|/2}),\ \ \ \ \bar{\phi}^{\prime}(x)=K+O(e^{-\delta|x|/2}),\quad x\to-\infty,\end{split}$
with
$\begin{split}c=\frac{1}{N}\log\left[\frac{\binom{m_{1}+2n_{1}}{m_{1}}}{\binom{m_{2}+2n_{2}}{m_{2}}}\frac{N!}{M!}\right]-\log 2\quad\text{and}\quad\delta=\frac{1}{2}\min\{1,K\}.\end{split}$
In fact, writing $\bar{g}_{m_{k+1},n_{k+1}}(x)=\bar{g}_{m_{k},n_{k}}(\bar{\phi}_{k}(x))$, since in any case we have chosen $m_{k}=n_{k}=0$, $k=1,2,3$, for the two sequences $(m_{k})$ and $(n_{k})$, we have $\bar{g}_{m_{3},n_{3}}=[g_{m_{3},n_{3}}+1]/2$ and $\bar{g}_{m_{2},n_{2}}=g_{m_{2},n_{2}}$ and thus
(4.9) $\bar{g}_{m_{3},n_{3}}(x)=(\bar{g}_{m_{2},n_{2}}\circ\bar{\phi}_{2})(x).$
We will actually use the asymptotic behaviors in (4.8) for the case $M=N=1$. On the other hand, we always have $\bar{\phi}_{k}=\phi_{k}$ is an increasing diffeomorphism on $\mathbb{R}$ which appears in (2.1) and (3.18) for the pair $(m_{k},2n_{k})$ when $k\not=3$.
By the construction of $(m_{k})$ and $(n_{k})$, we have $n_{k}\to\infty$ as $k\to\infty$ and thus the asymptotic behaviors of $\phi$ becomes more complicated. To study the asymptotic behaviors of $\bar{\phi}_{k}$ as $k\to\infty$, we need the following estimate
(4.10) $\begin{split}\frac{m_{k}}{2n_{k}+1}=O\left(\frac{1}{(\log(k+1))^{2}}\right),\quad k\to\infty\end{split}$
and thus
(4.11) $\begin{split}m_{k}+2n_{k}+1=\left(1+\frac{1}{(\log(k+1))^{2}}\right)(2n_{k}+1)\quad k\to\infty.\end{split}$
Recalling the notation $N_{k}=m_{k}+2n_{k}+1$, then we deduce from (4.1) that
(4.12) $\begin{split}\frac{N_{k+1}}{N_{k}}=1+O\left(\frac{1}{k}\right)\quad k\to\infty.\end{split}$
Moreover, denoting $\mathcal{N}_{k}=\sum_{j=1}^{k}N_{j}$, we have from (4.1) that
(4.13) $\begin{split}\mathcal{N}_{k}\sim(2\pi)^{\gamma-1}k^{\gamma},\quad k\to\infty.\end{split}$
The above estimates (4.10)–(4.13) are used in proving the lemmas below. We first have the following
Lemma 4.3 ([19]).
Let $m,n\in\mathbb{N}$ and put $N=m+2n+1$. Let $y>0$. Then
(4.14) $\begin{split}\log(h_{m,n}(y)-1)=&\ -\log\binom{m+2n}{m}N!+y+N\log y\\ &\ -2\log Q_{m}(y)-\log\left(1+\frac{y}{N}\right)+R(y,N),\end{split}$
where $R(y,N)\leq 4Bmy/[N(y+N-2m)]$ for all $y$ and all $N\geq N_{0}$ for some integer $N_{0}$ and some constant $B$.
We see that the error term $R(y,N)$ in (4.14) is uniformly bounded for all $y>0$ and all $N\geq N_{0}$ for an integer $N_{0}$. In particular, by (4.10) we see that $R(y,N)$ is of type $O((\log(k+1))^{-2})$ when $N\to\infty$. In terms of $\bar{g}_{m,n}$ defined in (4.6), equation in (4.14) takes the form
(4.15) $\begin{split}\log(\bar{g}_{m,n}(x)-1)=&\ -\log\binom{m+2n}{m}N!+e^{x}+Nx\\ &\ -2\log Q_{m}(e^{x})-\log\left(1+\frac{e^{x}}{N}\right)+R(e^{x},N)+\log 2\end{split}$
when $\bar{\chi}{m,n}(z)=(z+1)/2$, or takes the form
(4.16) $\begin{split}\log(\bar{g}_{m,n}(x)-1)=&\ -\log\binom{m+2n}{m}N!+e^{x}+Nx\\ &\ -2\log Q_{m}(e^{x})-\log\left(1+\frac{e^{x}}{N}\right)+R(e^{x},N)\end{split}$
when $\bar{\chi}_{m,n}(z)=z$. In the proof of the following lemmas, we shall always use the two estimates in (4.15) and (4.16) to consider $\bar{g}_{m,n}$ for different $\bar{\chi}_{m,n}$ respectively.
Also, for technical reasons, we shall not work with $\bar{g}_{m,n}$ directly. Since the function $\bar{g}_{m,n}:\mathbb{R}\to(1,\infty)$ is an increasing homeomorphism, there exists $\bar{s}_{m,n}\in\mathbb{R}$ such that
(4.17) $\begin{split}\bar{g}_{m,n}(\bar{s}_{m,n})=2.\end{split}$
For each pair of $(m,2n)$ of two nonnegative integers $m$ and $n$, we shall work with the function $\bar{g}_{m,n}(z+\bar{s}_{m,n})$ instead of the function $\bar{g}_{m,n}(z)$ directly. Now, for the function $\bar{g}_{m,n}(z)$, we may use (4.15) or (4.16) and slightly modify the proof of [19, Lemma 2.4] to prove the following
Lemma 4.4.
The constant $\bar{s}_{m,n}$ in (4.17) satisfies $\bar{s}_{m,n}=\log N+O(1)$ as $k\to\infty$.
For the two positive integers $M,N$ such that $M>N$, in the case (I) we shall assume that there exists a constant $C>1$ such that $\mathfrak{m}+2\mathfrak{n}\leq C(m+2n)$. Then clearly
(4.18) $\begin{split}M\leq CN.\end{split}$
The estimate in (4.12) implies that $C\to 1$ as $N\to\infty$. Below we begin to analyse the asymptotic behaviors of $\bar{\phi}_{k}(z)$ defined in (4.7). For the choice of $(m_{k})$ and $(n_{k})$ in the case (I) we know that $\bar{\phi}_{1}=\phi_{1}$ is the identity in the plane and the asymptotic behaviors of $\bar{\phi}_{2}$ has been given in (4.8), so below we only consider the case $k\geq 3$. Then, by slightly modifying the proof of [19, Lemma 2.5], we have the following
Lemma 4.5.
Let $M,N$ be two integers as before. Then
  • (1)
    if $\mathfrak{m}>m$, then $\bar{\phi}(x)$ has at least one fixed point $p$ and each fixed point $p$ of $\phi$ satisfies $p=\log N+O(1)$ as $k\to\infty$;
  • (2)
    if $\mathfrak{m}=m$, then $\bar{\phi}(x)$ has uniquely one fixed point $p$ such that $p=\log N+O(1)$ as $k\to\infty$;
  • (3)
    if $\mathfrak{m}<m$, then $\bar{\phi}(x)$ has uniquely one fixed point $p$ such that $p=\frac{l-1}{l}\log N+O(1)$ for an integer $l=M-N$ as $k\to\infty$;
  • (4)
    in all of the above three cases, $\bar{\phi}(x)<x$ for $x<p+O(1)$ and $\phi(x)>x$ for $x>p+O(1)$ and, moreover, we have $\bar{\phi}(x)\leq\log N+O(1)$ for all $x\leq\log N$.
Below we shall always assume that the terms $O(1)$ appearing are only dependent on the constant $C$, not on other variables. By the estimates in (4.15) or (4.16), we may slightly modify the proof of [19, Lemma 2.6] to obtain the following
Lemma 4.6.
For $\mathfrak{m},\mathfrak{n},m,n,M,N$ and $\phi:\mathbb{R}\to\mathbb{R}$ there exist positive constants $c_{1},\cdots,c_{8}$ depending only on the constant $C$ in (4.18) such that
$\begin{split}|\bar{\phi}(x)-x|\leq c_{1}e^{-x/2}\leq\frac{c_{1}}{N}\quad\text{for}\quad x>8\log N\end{split}$
and
$\begin{split}\left|\bar{\phi}(x)-\frac{M}{N}x+\frac{1}{N}\log\frac{M!}{N!}\right|\leq c_{2}e^{x}\quad\text{for}\quad x<\log N.\end{split}$
Moreover, with $\bar{s}_{m,n}$ defined in (4.17) we have
$\begin{split}\left|\bar{\phi}(x)-x\right|\leq c_{3}\quad\text{for}\quad x>\bar{s}_{m,n}\end{split}$
and
$\begin{split}\left|\bar{\phi}(x)-\frac{M}{N}x+\frac{1}{N}\log\frac{M!}{N!}\right|\leq c_{4}\quad\text{for}\quad x<\log N.\end{split}$
Finally,
$\begin{split}|\bar{\phi}^{\prime}(x)-1|\leq c_{5}\left(e^{-x/2}+\frac{1}{(\log(x+2\pi))^{2}}\right)\leq\frac{2c_{5}}{(\log(x+2\pi))^{2}}\quad\text{for}\quad x>8\log N\end{split}$
and
$\begin{split}\left|\bar{\phi}^{\prime}(x)-\frac{M}{N}\right|\leq\frac{c_{6}e^{x}}{N}\quad\text{for}\quad x<\log N,\end{split}$
as well as
$\begin{split}c_{7}\leq|\bar{\phi}^{\prime}(x)|\leq c_{8}\quad\text{for all}\quad x\in\mathbb{R}.\end{split}$
4.2. Definition of a quasiregular map
We begin to define a quasiregular map using the functions $\bar{g}_{m,n}$ defined in (4.6). By modifying these functions slightly, we obtain closely related functions $u_{m,n}$ and $v_{m,n}$ and then glue restrictions of these maps to half-strips along horizontal lines to obtain quasiregular maps $U$ and $V$ which are defined in the right and left half-planes respectively. Then we will glue these functions along the imaginary axis to obtain a quasiregular map $G$ in the plane.
Let $(m_{k})$ and $(n_{k})$ be the two sequences from Lemma 4.1 and Lemma 4.2 respectively. Then we give explicit expression for $\bar{g}_{m,n}$ defined in (4.6). On each strip
$\hat{\Pi}_{k}=\{z=x+iy:\quad 2\pi(k-1)\leq y\leq 2\pi k\}$
of the upper half-plane we will consider the functions
(4.19) $\bar{g}_{m_{k},n_{k}}(z)=g_{m_{k},n_{k}}\left(x+i\frac{y-2\pi N_{k-1}}{k}\right)$
when $k=1,2$ and
(4.20) $\bar{g}_{m_{k},n_{k}}(z)=\frac{1}{2}\left[g_{m_{k},n_{k}}\left(x+i\frac{y-2\pi N_{k-1}}{k}\right)+1\right]$
when $k\geq 3$ and on each strip
$\check{\Pi}_{-k}=\{z=x+iy:\quad 2\pi(-k)\leq y\leq 2\pi(-k+1)\}$
of the lower half-plane we will consider the functions
(4.21) $\bar{g}_{m_{k},n_{k}}(z)=g_{m_{k},n_{k}}\left(x-i\frac{y-2\pi N_{1-k}}{k}\right).$
In view of the explicit expressions of $\bar{g}_{m_{k},n_{k}}$ defined in (4.19), (4.20) and (4.21), the maps $U$, $V$ and $G$ will not commute with the complex conjugation. However, we shall only define them in the upper half-plane since the definitions in the lower half-plane are similar.
We begin by constructing the map $U$. Instead of $\bar{g}_{m,n}$ defined in (4.6), we consider the map
$\begin{split}u_{m,n}:\{z\in\mathbb{C}:\text{Re}\ z\geq 0\}\to\mathbb{C},\quad u_{m,n}(z)=\bar{g}_{m,n}(z+\bar{s}_{m,n}).\end{split}$
Note that $\bar{g}_{m,n}$ is increasing on the real line and maps $[0,\infty)$ onto $[2,\infty)$. Put $U(z)=u_{m_{k},n_{k}}(z)$ in the half-strip
$\begin{split}\hat{\Pi}_{k}^{+}=\{x+iy:x>0,\quad 2\pi(k-1)<y<2\pi k\}.\end{split}$
The function $U$ will be discontinuous along the horizontal lines $L^{+}_{k-1}=\{z=x+iy:x>0,y=2\pi(k-1)\}$ and $L^{+}_{k}=\{z=x+iy:x>0,y=2\pi k\}$. In order to obtain a continuous function we consider the function $\bar{\psi}_{k}:[0,\infty)\to[0,\infty)$ defined by $u_{m_{k+1},n_{k+1}}(x)=u_{m_{k},n_{k}}(\bar{\psi}_{k}(x))$. The function $\bar{\psi}_{k}$ is closely related to the function $\bar{\phi}$ considered in Lemma 4.5 and Lemma 4.6 and also the function $\bar{\phi}_{2}$ defined in (4.9). In fact, denoting by $\bar{\phi}_{k}$ the function $\bar{\phi}$ corresponding to $(m_{k},2n_{k})$ and $(m_{k+1},2n_{k+1})$, we have
$\begin{split}\bar{\psi}_{k}(x)=\bar{\phi}_{k}(x+\bar{s}_{m_{k+1},n_{k+1}})-\bar{s}_{m_{k},n_{k}}.\end{split}$
We then define $U:\{z\in\mathbb{C}:\text{Re}\ z\geq 0\}\to\mathbb{C}$ by interpolating between $u_{m_{k+1},n_{k+1}}$ and $u_{m_{k},n_{k}}$ as follows: if $2\pi(k-1)\leq y<2\pi k$, say $y=2\pi(k-1)+2\pi t$ where $10\leq t<$ 1, then we put
$\begin{split}U(x+iy)=u_{m_{k},n_{k}}((1-t)x+t\bar{\psi}_{k}(x)+iy)=u_{m_{k},n_{k}}(x+iy+t(\bar{\psi}_{k}(x)-x)).\end{split}$
We now define a function $V$ in the left half-plane. In order to do so, we define
$\begin{split}v_{m,n}:\{z\in\mathbb{C}:\text{Re}\ z<0\}\to\mathbb{C},\quad v_{m,n}(z)=\bar{g}_{m,n}\left(\frac{z}{m+2n+1}+\bar{s}_{m,n}\right).\end{split}$
Note that $v_{m,n}$ maps $(-\infty,0]$ monotonically onto $(1,2]$. Let $(m_{k})$ and $(n_{k})$ be as before and write $N_{k}=m_{k}+2n_{k}+1$. Put $\mathcal{N}_{k}=\sum_{j=1}^{k}N_{j}$. This time we would like to define $V(z)=v_{m_{k},n_{k}}(z)$ in the half-strip
[…]
but again this function would be discontinuous along the horizontal lines $\mathcal{L}^{-}_{k-1}=\{z=x+iy:x<0,y=2\pi\mathcal{N}_{k-1}\}$ and $\mathcal{L}^{-}_{k}=\{z=x+iy:x<0,y=2\pi\mathcal{N}_{k}\}$. In order to obtain a continuous function we again interpolate between $v_{m_{k+1},n_{k+1}}$ and $v_{m_{k},n_{k}}$. Similarly as before we consider the map $\bar{\psi}_{k}:(-\infty,0]\to(-\infty,0]$ defined by $v_{m_{k+1},n_{k+1}}(x)=v_{m_{k},n_{k}}(\bar{\psi}_{k}(x))$. Denoting by $\bar{\phi}_{k}$ the function $\bar{\phi}$ corresponding to $(m_{k},n_{k})$ and $(m_{k+1},n_{k+1})$, we have
$\begin{split}\bar{\psi}_{k}(x)=N_{k}\bar{\phi}_{k}\left(\frac{x}{N_{k+1}}+\bar{s}_{m_{k+1},n_{k+1}}\right)-N_{k}\bar{s}_{m_{k},n_{k}}.\end{split}$
Then we define $V:\{z\in\mathbb{C}:\text{Re}\ z\leq 0\}\to\mathbb{C}$ by interpolating between $v_{m_{k+1},n_{k+1}}$ and $v_{m_{k},n_{k}}$ as follows: if $2\pi\mathcal{N}_{k-1}\leq y<2\pi\mathcal{N}_{k}$, say $y=2\pi\mathcal{N}_{k-1}+2\pi N_{k}t$ where $10\leq t<$ 1, then we put
$\begin{split}V(x+iy)=v_{m_{k},n_{k}}((1-t)x+t\bar{\psi}_{k}(x)+iy)=v_{m_{k},n_{k}}(x+iy+t(\bar{\psi}_{k}(x)-x)).\end{split}$
This map $V$ is continuous along the horizontal lines $\mathcal{L}^{-}_{k}$ for any $k$ in the left half-plane.
Now we define our map $G$ by gluing $U$ and $V$ along the imaginary axis. In order to do this we note that by construction of $(m_{k})$ and $(n_{k})$ in Lemma 4.1 and Lemma 4.2 we have $U(iy)=V(ih(y))$ and thus $U(ig(y))=V(iy^{\gamma})$ for $y\geq 0$ with the maps $h$ and $g$ from Lemma 4.1. Therefore, we will consider a homeomorphism $Q$ of the right half-plane
$\begin{split}H^{+}=\{z\in\mathbb{C}:\text{Re}\ z\geq 0\}\end{split}$
onto itself, satisfying $Q(\bar{z})=\overline{Q(z)}$, such that $Q(\pm iy)=\pm ig(y)$ for $y\geq 0$ while $Q(z)=z$ for $\text{Re}\ z\geq 1$. We thus have to define $Q(z)$ in the strip $\{z\in\mathbb{C}:0<\text{Re}\ z<1\}$. For $|\text{Im}\ z|\geq 1$ we define $Q$ in this strip by interpolation; that is, we put
$\begin{split}Q(x\pm iy)=x\pm i((1-x)g(y)+xy)\quad\text{if}\quad 0<x<1\quad\text{and}\quad y\geq 1.\end{split}$
In the remaining part of the strip we define $Q$ by
$\begin{split}Q(z)=\left\{\begin{array}[]{ll}z|z|^{\gamma-1},&\text{if}\quad 0<|z|<1;\\ z,&\text{if}\quad|z|>1,\quad\text{but}\quad 0<\text{Re}\ z<1\quad\text{and}\quad 0\leq|\text{Im}\ z|\leq 1.\end{array}\right.\end{split}$
Note that $h(y)=y$ for $0\leq y\leq 2\pi$ since $m_{1}=n_{1}=0$ implies that $N_{1}=1$, and thus $g(y)=y^{\gamma}$ for $10\leq y<$ 1. Thus $Q(\pm iy)=\pm ig(y)$ also for $|y|\leq 1$. Moreover, we have $g(1)=1$, meaning that the above expressions for $Q$ do indeed coincide for $|\text{Im}\ z|=1$. We conclude that the map $W=U\circ Q$ satisfies
$\begin{split}W(\pm iy)=U(\pm ig(y))=V(\pm iy^{\gamma})\quad\text{for}\quad y\geq 0.\end{split}$
Let now $\rho\in(1/2,1)$. We choose the constant $\gamma=1/(2\rho-1)$ and put $\sigma=\rho/(2\rho-1)$. The hypothesis that $11/2<\rho<$ 1 corresponds to $\gamma>1$ as well as $\sigma>1$. Then we define the map
(4.22) $\begin{split}G(z)=\left\{\begin{array}[]{ll}W(z^{\rho}),&\text{if}\quad|\arg z|\leq\frac{\pi}{2\rho};\\ V(-(-z)^{\sigma}),&\text{if}\quad|\arg(-z)|\leq\frac{\pi}{2\sigma},\end{array}\right.\end{split}$
which is continuous on the imaginary axis and on the horizontal lines $L^{+}_{k}$ in the right half-plane and $\mathcal{L}^{-}_{k}$ in the left half-plane for any integer $k$ respectively. Here, for $\mu>0$, we denote by $z^{\mu}$ the principal branch of the power which is defined in $\mathbb{C}\setminus(-\infty,0]$.
4.3. Completion of the proof
With the definition of $G$ in (4.22) in previous section, here we first prove Theorem 1.6 under the assumption that $\rho\in(1/2,1)$. Now, we can follow exactly the same process as that in [5, Section 3.3] to estimate the dilation of $G$ and thus prove that $G$ is indeed quasiregular in the plane and also that $G$ satisfies the hypothesis of the Teichmüller–Wittich–Belinskii theorem. This theorem, together with the existence theorem for quasiconformal mappings, yields that there exists a quasiconformal homeomorphism $\tau:\mathbb{C}\to\mathbb{C}$ and a meromorphic function $F$ such that
(4.23) $\begin{split}G(z)=F(\tau(z))\quad\text{and}\quad\tau(z)\sim z\quad\text{as}\quad z\to\infty.\end{split}$
Then $F=G\circ\tau^{-1}$ is a locally univalent meromorphic function. Define $E=F/F^{\prime}$. Then $E$ is a Bank–Laine function. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A$.
Then we use similar arguments as in the proof of Theorem 1.3 to estimate the numbers of poles and zeros of $F$ respectively and also to determine the orders of $E$ and $A$. Again, two positive quantities $S_{1}(r)$ and $S_{2}(r)$ are written as $S_{1}(r)\asymp S_{2}(r)$ if there are two positive constants $d_{1}$ and $d_{2}$ such that $d_{2}S_{2}(r)\leq S_{1}(r)\leq d_{1}S_{2}(r)$ for all large $r$.
We first estimate the number of poles of $F$. Let $r>0$ and choose $k\in\mathbb{N}$ such that $2\pi(k-1)<r\leq 2\pi k$. We note from Lemma 4.1 that $\sum_{i=1}^{k}m_{i}\sim(2\pi)^{\delta\gamma-1}k^{\delta\gamma}$ when $0\leq\delta\gamma\leq 1$ and $2\sum_{i=1}^{k}m_{i}\sim(2\pi)^{\delta\gamma-1}(\log(k+2\pi))^{-2}k^{\delta\gamma}$ when $\delta\gamma>1$. It follows from the construction of $(m_{k})$ that the number of poles of $U$ satisfies
(4.24) $\begin{split}\sum_{i=1}^{k}m_{i}\leq n(r,\infty,U)\leq 2\sum_{i=1}^{k}m_{i}.\end{split}$
Thus the double inequalities in (4.24) yield $n(r,\infty,U)\asymp r^{\delta\gamma}$ when $0\leq\delta\gamma\leq 1$ and $n(r,\infty,U)\asymp(\log(r+2\pi))^{-2}r^{\delta\gamma}$ when $\delta\gamma>1$. Similarly, if $2\pi\mathcal{N}_{k-1}<r\leq 2\pi\mathcal{N}_{k}$, then the number of poles of $V$ satisfies
(4.25) $\begin{split}\sum_{i=1}^{k}m_{i}\leq n(r,\infty,V)\leq 2\sum_{i=1}^{k}m_{i}.\end{split}$
Since $\mathcal{N}_{k}\sim(2\pi)^{\gamma-1}k^{\gamma}$ as $k\to\infty$ by (4.13), then the double inequalities in (4.25) imply that $n(r,\infty,V)\asymp r^{\delta}$ when $0\leq\delta\gamma\leq 1$ and $n(r,\infty,V)\asymp(\log(r+2\pi))^{-2}r^{\delta}$ when $\delta\gamma>1$. Denote $\lambda=\delta\rho\gamma$. Since $\sigma=\rho\gamma$, we conclude that the number of poles of $G$ satisfies $n(r,\infty,G)\asymp r^{\lambda}$ when $0\leq\delta\gamma\leq 1$ and $n(r,\infty,G)\asymp(\log(r+2\pi))^{-2}r^{\lambda}$ when $\delta\gamma>1$. Now the relations in (4.23) imply that the number of poles of $F$ satisfies $n(r,\infty,F)\asymp r^{\lambda}$ when $0\leq\delta\gamma\leq 1$ and $n(r,\infty,F)\asymp(\log(r+2\pi))^{-2}r^{\lambda}$ when $\delta\gamma>1$. Thus $\lambda(f_{1})=\lambda$.
Then we estimate the number of zeros of $F$. To this end, we first estimate the number of zeros of $g_{m,n}+1$ for a fixed pair $(m,2n)$ of two nonnegative integers $m$ and $n$. Recall from [10, p. 7] that the Nevanlinna characteristic of the function $\exp(e^{z})$ satisfies
(4.26) $\begin{split}T(r,\exp(e^{z}))\sim\frac{e^{r}}{(2\pi^{3}r)^{1/2}},\quad r\to\infty.\end{split}$
By definition we see that the coefficients $A_{i}$ and $B_{j}$ in $g_{m,n}$ in (1.5) both satisfy $\sum_{i=0}^{\infty}A_{i}<\infty$ and $\sum_{j=0}^{\infty}|B_{j}|<\infty$ and also that $\sum_{i=0}^{\infty}iA_{i}<\infty$ and $\sum_{j=0}^{\infty}|jB_{j}|<\infty$. Write $g_{m,n}$ in (1.5) as $g_{m,n}(z)=\mathcal{R}(z)\exp(e^{z})$ for a function $\mathcal{R}(z)=R_{m,n}(e^{z})$. Then the first main theorem implies that
(4.27) $\begin{split}T(r,g_{m,n}+1)&=T(r,g_{m,n})+O(1)\\ &=T(r,\exp(e^{z}))+O(T(r,e^{mz}))+O(T(r,e^{nz}))+O(1).\end{split}$
On the other hand, we have the identity
$\begin{split}\frac{1}{g_{m,n}(z)+1}=1-\frac{1}{\mathcal{R}^{\prime}/\mathcal{R}+e^{z}}\frac{(g_{m,n}(z)+1)^{\prime}}{g_{m,n}(z)+1}.\end{split}$
Together with this identity, the lemma on the logarithmic derivative and the first main theorem imply that
(4.28) $\begin{split}m\left(r,\frac{1}{g_{m,n}+1}\right)&=O(\log rT(2r,g_{m,n}))+O\left(T(r,\mathcal{R}^{\prime}/\mathcal{R}+e^{z})\right)+O(1)\\ &=O(\log rT(2r,g_{m,n}))+O((m+2n+1)T(r,e^{z}))+O(1).\end{split}$
Therefore, for the two sequences $(m_{k})$ in Lemma 4.1 and $(n_{k})$ in Lemma 4.2, if $k=O(r)$ or $k^{\gamma}=O(r)$, then by (4.26) and (4.27) we have $\log T(r,g_{m,n}+1)\asymp r$ uniformly for all $m$ and $n$. Together with (4.28), the first main theorem implies that $\log N(r,0,g_{m,n}+1)\asymp r$ uniformly for all $m$ and $n$ and thus $\log n(r,0,g_{m,n}+1)\asymp r$ uniformly for all $m$ and $n$. Now, let $r>0$ and choose $k\in\mathbb{N}$ such that $2\pi(k-1)<r\leq 2\pi k$. Denote by $\hat{n}(r,0,\bar{g}_{m_{k},n_{k}})$ the number of zeros of $\bar{g}_{m_{k},n_{k}})$ in the first quadrant of the plane. Then, for each $k$, there are two integers $I=I(k)$ and $J=J(k)$ such that $0\leq I,J\leq k$ and
(4.29) $\begin{split}\hat{n}(r,0,\bar{g}_{m_{J},n_{J}})\leq\hat{n}(r,0,\bar{g}_{m_{i},n_{i}})\leq\hat{n}(r,0,\bar{g}_{m_{I},n_{I}})\end{split}$
for all $0\leq i\leq k$. Denote $u_{m_{I},n_{I}}=\bar{g}_{m_{I},n_{I}}(z+\bar{s}_{m_{I},n_{I}})$ and $u_{m_{J},n_{J}}=\bar{g}_{m_{J},n_{J}}(z+\bar{s}_{m_{J},n_{J}})$. It follows that the number of zeros of $U$ in the upper half-plane satisfies
(4.30) $\begin{split}\hat{n}(r,0,u_{m_{J},n_{J}})\leq\hat{n}(r,0,U)\leq\hat{n}(r,0,u_{m_{I},n_{I}}).\end{split}$
By the construction of the two sequences $(m_{k})$ in Lemma 4.1 and $(n_{k})$ in Lemma 4.2, it is easy to see that $\log\hat{n}(r,0,u_{m_{I},n_{I}})\asymp r$ and $\log\hat{n}(r,0,u_{m_{J},n_{J}})\asymp r$ uniformly for all $I$ and $J$. Then the double inequalities in (4.30) yield that $\log\hat{n}(r,0,U)\asymp r$. On the other hand, we choose $k\in\mathbb{N}$ such that $2\pi\mathcal{N}_{k-1}<r\leq 2\pi\mathcal{N}_{k}$. With the definitions of $\bar{g}_{m_{I},n_{I}}$ and $\bar{g}_{m_{I},n_{I}}$ in (4.29), now we have $v_{m_{I},n_{I}}=\bar{g}_{m_{I},n_{I}}(z/N_{I}+\bar{s}_{m_{I},n_{I}})$ and $v_{m_{J},n_{J}}=\bar{g}_{m_{J},n_{J}}(z/N_{J}+\bar{s}_{m_{J},n_{J}})$. It follows that the number of zeros of $V$ in the upper half-plane satisfies
(4.31) $\begin{split}\hat{n}(r,0,v_{m_{J},n_{J}})\leq\hat{n}(r,0,V)\leq\hat{n}(r,0,v_{m_{I},n_{I}}).\end{split}$
Since $N_{k}\sim\gamma(2\pi k)^{\gamma-1}$ as $k\to\infty$ by (4.1) and $\mathcal{N}_{k}\sim(2\pi)^{\gamma-1}k^{\gamma}$ as $k\to\infty$ by (4.13), then, by the construction of the two sequences $(m_{k})$ and $(n_{k})$, we have $\log\hat{n}(r,0,u_{m_{I},n_{I}})\asymp r^{1/\gamma}$ and $\log\hat{n}(r,0,u_{m_{J},n_{J}})\asymp r^{1/\gamma}$ for all $I$ and $J$. Then the double inequalities in (4.31) imply that $\log\hat{n}(r,0,V)\asymp r^{1/\gamma}$. Since $\sigma=\rho\gamma$, we conclude that the number of zeros of $G$ in the upper half-plane satisfies $\log\hat{n}(r,0,G)\asymp r^{\rho}$. Now the relations in (4.23) imply that the number of zeros of $F$ in the upper half-plane satisfies $\log\hat{n}(r,0,F)\asymp r^{\rho}$. Similarly, we may estimate the number of zeros of $F$ in the lower half-plane. Let $r>0$ and choose $k\in\mathbb{N}$ such that $2\pi(k-1)<r\leq 2\pi k$. It follows from the construction of $(n_{k})$ that the number of zeros of $U$ in the lower half-plane satisfies
(4.32) $\begin{split}\sum_{j=1}^{k}n_{j}\leq\check{n}(r,0,U)\leq 2\sum_{j=1}^{k}n_{j}.\end{split}$
Since $\mathcal{N}_{k}\sim(2\pi)^{\gamma-1}k^{\gamma}$ as $k\to\infty$ by (4.13) and since $2\sum_{j=1}^{k}n_{j}\sim\mathcal{N}_{k}$ as $k\to\infty$ by (4.11), the double inequalities in (4.32) yield $\check{n}(r,0,U)\asymp r^{\gamma}$. On the other, if we choose $k\in\mathbb{N}$ such that $2\pi\mathcal{N}_{k-1}<r\leq 2\pi\mathcal{N}_{k}$, then the number of zeros of $V$ in the lower half-plane satisfies
(4.33) $\begin{split}\sum_{j=1}^{k}n_{j}\leq\check{n}(r,0,V)\leq 2\sum_{j=1}^{k}n_{j}.\end{split}$
Since $\mathcal{N}_{k}\sim(2\pi)^{\gamma-1}k^{\gamma}$ as $k\to\infty$ by (4.13) and since $2\sum_{j=1}^{k}n_{j}\sim\mathcal{N}_{k}$ as $k\to\infty$ by (4.11), the double inequalities in (4.33) yield that $\check{n}(r,0,V)\asymp r$. Since $\sigma=\rho\gamma$, we conclude that the number of zeros of $G$ in the lower half-plane satisfies $\check{n}(r,0,G)\asymp r^{\rho\gamma}$. Now the relations in (4.23) imply that the number of zeros of $F$ in the lower half-plane satisfies $\check{n}(r,0,F)\asymp r^{\rho\gamma}$. By combining the two estimates for $\hat{n}(r,0,F)$ and $\check{n}(r,0,F)$ together, we conclude that $\log n(r,0,F)\asymp r^{\rho}$ and thus $\lambda(f_{2})=\infty$.
Now we show that $\rho(A)=\rho$ by using similar arguments as in Subsection 3.4. Now, with the choice of the two sequences $(m_{k})$ and $(n_{k})$ in Lemma 4.1 and Lemma 4.2 and the explicit expressions of $\bar{g}_{m_{k},n_{k}}$ in (4.19), (4.20) and (4.21), we may also suppose that the two estimates in (3.43) and (3.44) hold for $\bar{g}_{m_{k},n_{k}}$ for all pairs $(m_{k},2n_{k})$. Clearly, with the notation $L=m+2n$ and $L_{k}=m_{k}+2n_{k}$, this implies that there is a set $\Omega_{2}$ of finite linear measure such that
$\left\{\begin{array}[]{ll}|v_{m,n}(z)|\leq\exp((L+2\varepsilon)r/N)\left|\exp(e^{z/N})\right|&\text{for}\quad z\in H^{-}\quad\text{and}\quad|z|\not\in\Omega_{2},\\ |u_{m,n}(z)|\leq\exp((L+2\varepsilon)r)\left|\exp(e^{z})\right|&\text{for}\quad z\in H^{+}\quad\text{and}\quad|z|\not\in\Omega_{2}\end{array}\right.$
for all pairs $(m,2n)$. Let $x+iy\in H^{-}$ with $\text{Im}\ z\geq 0$ and $k\in\mathbb{N}$ with $2\pi\mathcal{N}_{k-1}\leq\text{Im}z<2\pi\mathcal{N}_{k}$. Then, with $t=(y-2\pi\mathcal{N}_{k-1})/{2\pi N_{k}}$ where $10\leq t<$ 1, we have a similar estimate as in (3.46), which implies that
(4.34) $\begin{split}|V(z)|\leq\exp((L_{k}+3\varepsilon)r)\end{split}$
for $z\in H^{-}$ and $|z|\not\in\Omega_{2}$. On the other hand, for $z\in H^{+}$ and $|z|\not\in\Omega_{2}$ we have
(4.35) $\begin{split}|\bar{g}_{m,n}(z)|\leq\exp((L+2\varepsilon)r)\exp\left(e^{\text{Re}\ z}\right).\end{split}$
Again, assuming that $\text{Im}\ z\geq 0$, we choose $k\in\mathbb{N}$ such that $2\pi(k-1)\leq\text{Im}\ z<2\pi k$. Then, with $U(z)=u_{m_{k},n_{k}}(q(z))=\bar{g}_{m_{k},n_{k}}(q(z)+\bar{s}_{m_{k},n_{k}})$ where $q(z)$ is defined by $q(x+iy)=x+iy+t(\bar{\psi}_{k}(x)-x)$, we have
(4.36) $\begin{split}|U(z)|\leq\exp((L_{k}+2\varepsilon)\text{Re}(q(z))+\bar{s}_{m_{k},n_{k}})\exp\left(e^{\text{Re}(q(z))+\bar{s}_{m_{k},n_{k}}}\right).\end{split}$
Noting that $N_{k}=O(k^{\gamma-1})=O(|z|^{\gamma-1})$ by the construction of $(m_{k})$ in Lemma 4.1 and $(n_{k})$ in Lemma 4.2, we have by Lemma 4.4 that
(4.37) $\begin{split}\bar{s}_{m_{k},n_{k}}=\log N_{k}+O(1)=O(\log|z|).\end{split}$
Then, since $L_{k}=N_{k}-1$, we deduce from (4.36) and (4.37) that
$\begin{split}\log|U(z)|\leq&\ (L_{k}+2\varepsilon)\text{Re}(q(z))+\bar{s}_{m_{k},n_{k}}+\exp(\text{Re}(q(z))+\bar{s}_{m_{k},n_{k}})\\ \leq&\ \exp((1+o(1))|z|)+O(|z|^{\gamma})\leq\exp((1+o(1))|z|),\quad z\to\infty\quad\text{in}\quad H^{+}.\end{split}$
Together with (4.34) we conclude that
$\begin{split}\log|G(z)|\leq\exp((1+o(1))|z|^{\rho})\end{split}$
as $|z|\to\infty$ outside a set $\Omega_{3}$ of finite linear measure and hence (4.23) yields that
$\begin{split}\log|F(z)|\leq\exp((1+o(1))|z|^{\rho})\end{split}$
as $|z|\to\infty$ outside a set $\Omega_{4}$ of finite linear measure. Together with (4.23), the lemma on the logarithmic derivative now implies that $E=F/F^{\prime}$ satisfies
$\begin{split}m\left(r,\frac{1}{E}\right)=O(r^{\rho})\end{split}$
outside a set of finite linear measure. Since all zeros of $E$ are simple, the previous estimates on the poles and zeros of $F$ imply that $\log N(r,0,E)\asymp r^{\rho}$. Then the above estimate together with the first main theorem imply that $\log T(r,E)\asymp r^{\rho}$. Then the lemma on the logarithmic derivative, together with (1.8), implies that
$\begin{split}m(r,A)=2m\left(r,\frac{1}{E}\right)+O(\log rT(r,E))=O(r^{\rho}).\end{split}$
Thus $\rho(A)\leq\rho$. On the other hand, we note that $g_{1,1}$ is analytic in $\hat{\Pi}_{1}^{+}$ as well as in $\check{\Pi}_{-1}^{+}$. Let $\bar{\eta}:H^{+}\to H^{+}$ be the map such that $\bar{\eta}=z^{\rho}$. Denote $\bar{Q}=Q+\bar{s}_{1,1}$. Note that $\bar{\eta}$ and $\bar{Q}$ both commute with the complex conjugation and also that are both analytic for $z$ in the right half-plane such that $\text{Re}\ z\geq 1$. Then, for $\tau$ in (4.23), by the same arguments as in the proof of Theorem 1.3 we may choose a curve in $(\tau\circ\bar{\eta}^{-1}\circ\bar{Q}^{-1})(\hat{\Pi}_{1}^{+}\cup\check{\Pi}_{-1}^{+})$, say $\mathcal{L}^{+}$, such that $(\bar{Q}\circ\bar{\eta}\circ\tau^{-1})(\mathcal{L}^{+})$ is the positive real axis and also that $\tau$ satisfies $|\tau^{\prime}(z)|\leq 2|z|$ for all $z\in(\bar{\eta}^{-1}\circ\bar{Q}^{-1})(\mathcal{L}^{+})$ such that $|z|$ is large. Then, by the relation in (1.8) as well as the definitions of $u_{m,n}$ in Subsection 4.2 and the relation $F=G\circ\tau^{-1}$, we easily see that $|A|\geq d_{3}|z|^{\rho}$ for some positive constant $d_{3}$ and all $z\in\mathcal{L}^{+}$ such that $|z|$ is large, implying that $\rho(A)\geq\rho$. Hence $\rho(A)=\rho$.
To complete the proof under the assumption $\rho\in(1/2,1)$, we show that the second order differential equation (1.1) with the constructed coefficient $A$ cannot have any other solution $f_{3}$ linearly independent from $f_{1}$ such that $\lambda(f_{3})<\infty$. Otherwise, there are two nonzero constants $d_{4}$ and $d_{5}$ such that $f_{3}=d_{4}f_{1}+d_{5}f_{2}$ and thus
(4.38) $\begin{split}f_{3}^{2}=(d_{4}f_{1}+d_{5}f_{2})^{2}=d_{5}^{2}f_{1}^{2}\left(F+d_{4}/d_{5}\right)^{2}=d_{5}^{2}f_{1}^{2}\left(G\circ\tau^{-1}+\frac{d_{4}}{d_{5}}\right)^{2},\end{split}$
where $\tau$ is defined in (4.23). By the explicit expressions of $\bar{g}_{m_{k},n_{k}}$ in (4.19), (4.20) and (4.21) we see that the zeros $\check{n}(r,0,f_{3})$ of $f_{3}$ in the lower half-plane satisfies $\log\check{n}(r,0,f_{3})\geq d_{6}r^{\rho}$ for some positive constant $d_{6}$ and all large $r$, which implies that $\lambda(f_{3})=\infty$, a contradiction. Thus the desired result follows.
Finally, recalling that $m_{k}=n_{k}=0$ for $k=1,2,3$ in the two sequences $(m_{k})$ and $(n_{k})$, we may follow exactly the same process as that in [5, Section 5] together with the comments in Subsection 3.5 to extend the above results to the case $\rho\in(n/2,n)$ for any positive integer $n$, except that the corresponding two functions $G_{0}$ and $G_{1}$ there are now defined by using the function $G$ in the case $\rho\in(1/2,1)$ here. Then we obtain a locally univalent meromorphic function $F$. The numbers of poles and zeros of $F$ can be estimated in a similar way as in the case $\rho\in(1/2,1)$ and the order of the corresponding coefficient $A$ of the second order differential equation (1.1) can also be determined by using similar arguments as there. Moreover, in certain region of the plane we may use similar arguments as for equation (4.38) to show that any other solution $f_{3}$ linearly independent from $f_{1}$ of the second order differential equation (1.1) with the constructed coefficient $A$ cannot satisfy $\lambda(f_{3})<\infty$. We shall not repeat the details. This completes the proof of Theorem 1.6.
5. Proof of Theorem 1.7
The proof for Theorem 1.7 will be similar as that in Section 3 and that in Section 4 in many parts. So we shall keep the details minimal in the following.
Also, we first prove Theorem 1.7 for the case $n=1$. We let $\lambda_{1}\in[0,1]$ and $\lambda_{2}\in[0,1]$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$. Then we shall choose two sequences $(m_{k})$ and $(n_{k})$ associated with $\lambda_{1}$ and $\lambda_{2}$ respectively in the same way as in Subsection 3.1, namely,
  • (I)
    if $\lambda_{1}\leq\lambda_{2}<1$, then we choose the sequence $(m_{k})$ and $(n_{k})$ constructed in Subsection 3.1 respectively; moreover, by slightly modifying the proof of Lemma 3.1, we may choose a different sequence $(n_{k})$ from the one there so that the function $\tilde{h}_{2}$ related to $\lambda_{2}$ can also satisfy $\tilde{h}_{2}(x)=(1+o(1))\lambda_{3}\log r$ for any given number $\lambda_{3}\in[0,\infty)$ in the case $\lambda_{2}=0$;
  • (II)
    if $\lambda_{1}<\lambda_{2}=1$, then we choose the sequence $(m_{k})$ constructed in Subsection 3.1 and $(n_{k})$ in the way that $n_{k}=0$ for $k=1,2,3$ and $n_{k}=1$ for all $k\geq 4$ respectively;
  • (III)
    if $\lambda_{1}=\lambda_{2}=1$, then we choose the sequence $(m_{k})$ and $(n_{k})$ in the way that $m_{k}=0$ for $k=1,2,3$ and $m_{k}=1$ for all $k\geq 4$ and $n_{k}=0$ for $k=1,2,3$ and $n_{k}=1$ for all $k\geq 4$ respectively.
With the definitions of $\bar{g}_{m,n}$ in (4.6), for each pair $(m,2n)$ of two nonnegative integers $m$ and $n$ and each pair $(\mathfrak{m},2\mathfrak{n})$ of two nonnegative integers $\mathfrak{m}$ and $\mathfrak{n}$, we shall consider the function $\bar{\phi}:\mathbb{R}\to\mathbb{R}$ defined in (4.7), i.e.,
$\begin{split}(\bar{\chi}_{\mathfrak{m},\mathfrak{n}}\circ g_{\mathfrak{m},\mathfrak{n}})(x)=\bar{g}_{\mathfrak{m},\mathfrak{n}}(x)=(\bar{g}_{m,n}\circ\bar{\phi})(x)=(\bar{\chi}_{m,n}\circ g_{m,n}\circ\bar{\phi})(x).\end{split}$
Put $N=m+2n+1$ and $M=\mathfrak{m}+2\mathfrak{n}+1$. We will consider the functions $\bar{\phi}$ for the case that $m=m_{k}$, $n=n_{k}$, $\mathfrak{m}=m_{k+1}$ and $\mathfrak{n}=n_{k+1}$, that is, $M=N_{k+1}$ and $N=N_{k}$. Thus we will consider the behavior of $\bar{\phi}_{k}$ corresponding to the pair $(m_{k},2n_{k})$ of two nonnegative integers $m_{k}$ and $n_{k}$ as $k\to\infty$, but in order to simplify the formulas in most situations below we suppress the dependence of $\bar{\phi}$ from $m,n$ and $\mathfrak{m},\mathfrak{n}$ from the notation.
In all the three cases (I), (II) and (III), we have $N\leq 4$ in any case and thus the asymptotic behaviors of $\bar{\phi}$ described in Lemma 2.1 apply to all $\bar{\phi}_{k}$ in the case $\bar{\chi}_{m_{k+1},n_{k+1}}=\bar{\chi}_{m_{k},n_{k}}$ and the asymptotic behaviors of $\bar{\phi}$ described in (4.8) apply to all $\bar{\phi}_{k}$ in the case $\bar{\chi}_{m_{k+1},n_{k+1}}(z)=(z+1)/2$ and $\bar{\chi}_{m_{k},n_{k}}(z)=z$. Moreover, the asymptotic behaviors of $\bar{\phi}$ for the case $\bar{\chi}_{m_{k+1},n_{k+1}}=z$ and $\bar{\chi}_{m_{k},n_{k}}(z)=(z+1)/2$ can be easily obtained as in (4.8). We omit presenting the details. Also, since $N\leq 4$ in any case, for the constant $\bar{s}_{m,n}\in\mathbb{R}$ such that $\bar{g}_{m,n}(\bar{s}_{m,n})=2$ defined in (4.17), we have $\bar{s}_{m,n}=O(1)$ for all $m,n$, which can be proved by slightly modifying the proof of Lemma 3.2.
Note that $m_{k}=0$ for $k=1,2,3$ and $n_{k}=0$ for $k=1,2,3$ for each two sequences $(m_{k})$ and $(n_{k})$ in (I), (II) and (III). We begin to define a quasiregular maps $U$, $V$ and $G$ by gluing the functions $\bar{g}_{m,n}$ defined in (4.6), but with the new sequences $(m_{k})$ and $(n_{k})$ chosen above. Also, we have $U(iy)=V(iy)$ on the imaginary axis. Again, we first work with the map
$\bar{g}_{m,n}\circ\chi_{m,n}=\bar{\chi}_{m,n}\circ g_{m,n}\circ\chi_{m,n}=\bar{\chi}_{m,n}\circ\tilde{g}_{m,n},$
where $\chi_{m,n}$ is defined in (3.18). The explicit expressions for the functions $\bar{g}_{m,n}\circ\chi_{m,n}$ with different pairs $(m_{k},2n_{k})$ of two nonnegative integers $m_{k}$ and $n_{k}$ are given as follows. With the definitions of $\tilde{g}_{m_{k},n_{k}}$ defined in (3.20), on each strip
$\begin{split}\hat{\mathbf{\Pi}}_{k}=\{x+iy:\quad y\geq 0,\quad 2\pi\mathcal{N}_{k-1}<y<2\pi\mathcal{N}_{k}\}\end{split}$
of the upper half-plane we have
(5.1) $\begin{split}(\bar{g}_{m_{k},n_{k}}\circ\chi_{m_{k},n_{k}})(z)=\left\{\begin{array}[]{ll}\frac{1}{2}\left[g_{m_{k},n_{k}}\left(\frac{x}{l}+i\frac{y-2\pi\mathcal{N}_{k-1}}{N_{k}}\right)+1\right],&x\geq 0,\\ \frac{1}{2}\left[g_{m_{k},n_{k}}\left(\frac{x+iy-2i\pi\mathcal{N}_{k-1}}{N_{k}}\right)+1\right],&x\leq 0\end{array}\right.\end{split}$
if $n_{k}=1$ and
(5.2) $\begin{split}(\bar{g}_{m_{k},n_{k}}\circ\chi_{m_{k},n_{k}})(z)=\left\{\begin{array}[]{ll}g_{m_{k},n_{k}}\left(\frac{x}{l}+i\frac{y-2\pi\mathcal{N}_{k-1}}{N_{k}}\right),&x\geq 0,\\ g_{m_{k},n_{k}}\left(\frac{x+iy-2i\pi\mathcal{N}_{k-1}}{N_{k}}\right),&x\leq 0\end{array}\right.\end{split}$
if $n_{k}=0$. On the other hand, on each strip
$\check{\mathbf{\Pi}}_{-k}=\{z=x+iy:\quad y\leq 0,\quad 2\pi\mathcal{N}_{-k}\leq y\leq 2\pi\mathcal{N}_{-k+1}\}$
of the lower half-plane we have
(5.3) $\begin{split}(\bar{g}_{m_{k},n_{k}}\circ\chi_{m_{k},n_{k}})(z)=\left\{\begin{array}[]{ll}g_{m_{k},n_{k}}\left(\frac{x}{l}-i\frac{y-2\pi\mathcal{N}_{1-k}}{N_{k}}\right),&x\geq 0,\\ g_{m_{k},n_{k}}\left(\frac{x-iy+2i\pi\mathcal{N}_{1-k}}{N_{k}}\right),&x\leq 0.\end{array}\right.\end{split}$
In the above new definitions, we shall choose $l=1$ in the case (I) and $l=3$ in the case (II) and $l=4$ in the case (III). Note that $n_{k}=0$ for $k=1,2,3$ and thus we have the expressions in (5.2) in $\hat{\mathbf{\Pi}}_{1}$, $\hat{\mathbf{\Pi}}_{2}$ and $\hat{\mathbf{\Pi}}_{3}$. Also note that the two expressions in each of the definition of $\bar{g}_{m_{k},n_{k}}\circ\chi_{m_{k},n_{k}}$ in (5.1), (5.2) or (5.3) coincide on the imaginary axis.
In view of these definitions, the maps $U$, $V$ and $G$ will not commute with the complex conjugation, but we shall only define them in the upper half-plane since the definitions in the lower half-plane are similar.
The definitions for $U$ and $V$ will be similar as that in Subsection 3.2, except that we shall use the explicit expressions of $\bar{g}_{m,n}\circ\chi_{m_{k},n_{k}}=\bar{\chi}_{m,n}\circ\tilde{g}_{m,n}$ defined in (5.1) and (5.2), instead of the functions $\tilde{g}_{m,n}$ defined in (3.20). To construct the map $U$, we define
$\begin{split}u_{m,n}:\{z\in\mathbb{C}:\text{Re}\ z\geq 0\}\to\mathbb{C},\quad u_{m,n}(z)=\bar{g}_{m,n}(\chi_{m,n}(z)+\bar{s}_{m,n}),\end{split}$
where $\chi_{m,n}$ is defined in (3.18). Note that $u_{m,n}$ is increasing on the real line and maps $[0,\infty)$ onto $[2,\infty)$. For each two chosen sequences $(m_{k})$ and $(n_{k})$, we write $N_{k}=m_{k}+2n_{k}+1$. Put $U(z)=u_{m_{k},n_{k}}$ in the half-strip
$\begin{split}\hat{\mathbf{\Pi}}_{k}^{+}=\{x+iy:x>0,\quad 2\pi\mathcal{N}_{k-1}<y<2\pi\mathcal{N}_{k}\}.\end{split}$
Then we consider the map
$\begin{split}u_{m_{k},n_{k}}:\{z\in\mathbb{C}:\text{Re}\ z\geq 0\}\to\mathbb{C},\quad u_{m_{k},n_{k}}(z)=\bar{g}_{m_{k},n_{k}}(\chi_{m_{k},n_{k}}(z)+\bar{s}_{m_{k},n_{k}})\end{split}$
when $n_{k}=0$ and
$\begin{split}u_{m_{k},n_{k}}:\{z\in\mathbb{C}:\text{Re}\ z\geq 0\}\to\mathbb{C},\quad u_{m_{k},n_{k}}(z)=\frac{\bar{g}_{m_{k},n_{k}}(\chi_{m_{k},n_{k}}(z)+\bar{s}_{m_{k},n_{k}})+1}{2}\end{split}$
when $n_{k}=1$. Consider the function $\bar{\psi}_{k}:[0,\infty)\to[0,\infty)$ defined by $u_{m_{k+1},n_{k+1}}(x)=u_{m_{k},n_{k}}(\bar{\psi}_{k}(x))$. Then $\bar{\psi}_{k}$ takes the form
$\begin{split}\bar{\psi}_{k}(x)=l\bar{\phi}_{k}\left(\frac{x}{l}+\bar{s}_{m_{k+1},n_{k+1}}\right)-l\bar{s}_{m_{k},n_{k}}.\end{split}$
Then we may interpolate in the same way as in Subsection 3.2 to obtain a map $U$ which is continuous along the horizontal lines $\mathcal{L}^{+}_{k}=\{z=x+iy,x>0,y=2\pi\mathcal{N}_{k}\}$ for any $k$ in the right half-plane. We omit those details. On the other hand, to construct a function $V$ in the left half-plane, we define
$\begin{split}v_{m,n}:\{z\in\mathbb{C}:\text{Re}\ z<0\}\to\mathbb{C},\quad v_{m,n}(z)=\bar{g}_{m,n}\left(\chi_{m,n}(z)+\bar{s}_{m,n}\right),\end{split}$
where $\chi_{m,n}$ is defined in (3.18). Note that $v_{m,n}$ maps $(-\infty,0]$ monotonically onto $(1,2]$. Let $(m_{k})$ and $(n_{k})$ be the two sequences as before. This time we would like to define $V(z)=v_{m_{k},n_{k}}(z)$ in the half-strip
[…]
Then we consider the map
$\begin{split}v_{m_{k},n_{k}}:\{z\in\mathbb{C}:\text{Re}\ z\leq 0\}\to\mathbb{C},\quad v_{m_{k},n_{k}}(z)=\bar{g}_{m_{k},n_{k}}(\chi_{m_{k},n_{k}}(z)+\bar{s}_{m_{k},n_{k}})\end{split}$
when $n_{k}=0$ and
$\begin{split}v_{m_{k},n_{k}}:\{z\in\mathbb{C}:\text{Re}\ z\leq 0\}\to\mathbb{C},\quad v_{m_{k},n_{k}}(z)=\frac{\bar{g}_{m_{k},n_{k}}(\chi_{m_{k},n_{k}}(z)+\bar{s}_{m_{k},n_{k}})+1}{2}\end{split}$
when $n_{k}=1$. Consider the map $\bar{\psi}_{k}:(-\infty,0]\to(-\infty,0]$ defined by $v_{m_{k+1},n_{k+1}}(x)=v_{m_{k},n_{k}}(\bar{\psi}_{k}(x))$. Then $\bar{\psi}_{k}$ takes the form
$\begin{split}\bar{\psi}_{k}(x)=N_{k}\bar{\phi}_{k}\left(\frac{x}{N_{k+1}}+\bar{s}_{m_{k+1},n_{k+1}}\right)-N_{k}\bar{s}_{m_{k},n_{k}}.\end{split}$
Then we may interpolate in the same way as in Subsection 3.2 to obtain a map $V$ which is continuous along the horizontal lines $\mathcal{L}^{-}_{k}=\{z=x+iy,x<0,y=2\pi\mathcal{N}_{k}\}$ for any $k$ in the left half-plane. We omit those details.
Note that $U$ and $V$ coincide on the imaginary axis. With the definitions of the maps $U$ and $V$, we define the map
(5.4) $\begin{split}G(z)=\left\{\begin{array}[]{ll}U(z),&\text{if}\quad|\arg z|\leq\frac{\pi}{2};\\ V(z),&\text{if}\quad|\arg(-z)|\leq\frac{\pi}{2},\end{array}\right.\end{split}$
which is continuous along the horizontal lines $\mathcal{L}^{+}_{k}$ in the right half-plane and $\mathcal{L}^{-}_{k}$ in the left half-plane for any integer $k$ respectively.
Then we can follow exactly the same process as in Subsection 3.3 to show that the map $G$ defined in (5.4) is indeed quasiregular in the plane and satisfies the hypothesis of the Teichmüller–Wittich–Belinskii. This theorem, together with the existence theorem for quasiconformal mappings, yields that there exists a quasiconformal homeomorphism $\tau:\mathbb{C}\to\mathbb{C}$ and a meromorphic function $F$ such that
(5.5) $\begin{split}G(z)=F(\tau(z))\quad\text{and}\quad\tau(z)\sim z\quad\text{as}\quad z\to\infty.\end{split}$
Then $F=G\circ\tau^{-1}$ is a locally univalent meromorphic function. Define $E=F/F^{\prime}$. Then $E$ is a Bank–Laine function. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A$.
We now estimate the numbers of poles and zeros of $F$ respectively. To estimate the number of poles of $F$, with the previous notation $S_{1}(r)\asymp S_{2}(r)$, we may use the same arguments as in Subsection 3.4 to obtain that $n(r,\infty,F)\asymp(\log r)^{2}$ when $\lambda_{1}=0$ and $n(r,\infty,F)\asymp r^{\lambda_{1}}$ when $0<\lambda_{1}\leq 1$. Thus $\lambda(f_{1})=\lambda_{1}$.
To estimate the number of zeros of $F$, we combine the methods in Subsection 3.4 and Subsection 4.3 together. Let $r>0$ and choose $k\in\mathbb{N}$ such that $2\pi\mathcal{N}_{k-1}<r\leq 2\pi\mathcal{N}_{k}$. Denote by $\hat{n}(r,0,\bar{g}_{m_{k},n_{k}})$ the number of zeros in the first quadrant of the plane. Then, for each $k$, there are two integers $I=I(k)$ and $J=J(k)$ such that $0\leq I,J\leq k$ and the double inequalities in (4.29) hold for all $0\leq i\leq k$, with the choice of the two sequences $(m_{k})$ and $(n_{k})$ now. For each $k$, we define $U_{I}$ in the way that $\bar{g}_{m_{i},n_{i}}$ with $n_{i}=1$ is replaced by $\bar{g}_{m_{I},n_{I}}$ in the right half-plane and $U_{J}$ in the way that $\bar{g}_{m_{i},n_{i}}$ with $n_{i}=1$ is replaced by $\bar{g}_{m_{J},n_{J}}$ in the left half-plane. Correspondingly, we have $u_{m_{I},n_{I}}=\bar{g}_{m_{I},n_{I}}(z/l+\bar{s}_{m_{I},n_{I}})$ and $u_{m_{J},n_{J}}=\bar{g}_{m_{J},n_{J}}(z/l+\bar{s}_{m_{J},n_{J}})$ and $v_{m_{I},n_{I}}=\bar{g}_{m_{I},n_{I}}(z/N_{I}+\bar{s}_{m_{I},n_{I}})$ and $v_{m_{J},n_{J}}=\bar{g}_{m_{J},n_{J}}(z/N_{J}+\bar{s}_{m_{J},n_{J}})$. It follows that the number of zeros of $U$ in the upper half-plane satisfies
(5.6) $\begin{split}\hat{n}(r,0,U_{J})\leq\hat{n}(r,0,U)\leq\hat{n}(r,0,U_{I})\end{split}$
and also that the number of zeros of $V$ in the upper half-plane satisfies
(5.7) $\begin{split}\hat{n}(r,0,V_{J})\leq\hat{n}(r,0,V)\leq\hat{n}(r,0,V_{I}).\end{split}$
Note that $\mathcal{N}_{k}\sim lk$ as $k\to\infty$ by (3.31). We recall that, in the case $10\leq\lambda_{1}\leq\lambda_{2}<$ 1, in the proof of Lemma 3.1 the subsequence $(m_{k_{i}})$ is chosen to satisfy the relation in (3.5) so that the double inequalities in (3.6) or (3.7) hold. For the two sequences $(m_{k})$ and $(n_{k})$ we choose, this implies that $\log\hat{n}(r,0,U_{I})\asymp j=j(r)$ for the greatest $j$ such that $n_{k_{j}}=1$ in the case $\lambda_{2}<1$ and $\log\hat{n}(r,0,U_{I})\asymp r$ in the case $\lambda_{2}=1$. It follows that $\hat{n}(r,0,U_{I})\asymp r^{\lambda_{3}}$ or $\log\hat{n}(r,0,U_{I})\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\log\hat{n}(r,0,U_{I})\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Such estimates also hold for $\log\hat{n}(r,0,U_{J})$, $\log\hat{n}(r,0,V_{I})$ and $\log\hat{n}(r,0,V_{J})$. Recall that $2\sum_{j=1}^{k}n_{j}\sim(2\pi)^{-1}(\log(2\pi k))^{2}$ as $k\to\infty$ when $\lambda_{2}=0$ and that $2\sum_{j=1}^{k}n_{j}\sim(2\pi)^{\lambda_{2}-1}k^{\lambda_{2}}$ as $k\to\infty$ when $0<\lambda_{2}\leq 1$, which are pointed out in Subsection 3.4. Then the double inequalities in (5.6) and (5.7) imply that the number of zeros of $G$ in the upper half-plane satisfies $\hat{n}(r,0,G)\asymp r^{\lambda_{3}}$ or $\log\hat{n}(r,0,G)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\log\hat{n}(r,0,G)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Now the relations in (5.5) imply that the number of zeros of $F$ in the upper half-plane satisfies $\hat{n}(r,0,G)\asymp r^{\lambda_{3}}$ or $\log\hat{n}(r,0,F)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\log\hat{n}(r,0,F)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Also, for the number of zeros of $F$ in the lower half-plane, we can use the same method as in Subsection 3.4 to show that $\check{n}(r,0,F)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\check{n}(r,0,F)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. By combining the two estimates for $\hat{n}(r,0,F)$ and $\check{n}(r,0,F)$ together, we conclude that $n(r,0,G)\asymp r^{\lambda_{3}}$ or $\log n(r,0,F)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\log n(r,0,F)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$ and thus $\lambda(f_{2})=\infty$.
Moreover, by similar arguments as in Subsection 3.4 we can estimate the modulus of $\log\log|G|$ and $\log\log|F|$ to conclude that the resulting Bank–Laine function $E$ satisfies $m(r,1/E)=O(r)$ and further that $\rho(A)\leq 1$. On the other hand, by considering the modulus of $E$ on the positive real axis we may conclude that $\rho(A)\geq 1$. Hence $\rho(A)=1$. Also, for any other solution $f_{3}$ linearly independent from $f_{1}$ of the second order differential equation (1.1) with the constructed coefficient $A$ we may consider the equation (4.38) and show that $f_{3}$ cannot satisfy $\lambda(f_{3})<\infty$. This completes the proof for the case $n=1$.
Finally, the above results can be extended to the case $\rho=n$ for any positive integer $n$ in the same way as that in [5, Section 5] together with the comments in Subsection 3.5 and also that in Subsection 4.3. We shall not repeat the details. This completes the proof of Theorem 1.7.
6. Acknowledgements
The author would like to greatly thank professor Walter Bergweiler for having enlightening discussions on the Bank–Laine conjecture and the quasiconformal surgery during his visit to the University of Kiel in July 2024. In particular, during the discussions the author specified the main obstacles in the proof of Theorem 1.4 and Theorem 1.6. Some ideas for overcoming these obstacles were proposed during the discussions.